### v1

# Tetrahedron Conjecture in the $\ell_{2}$-norm

Abstract
The famous Tetrahedron Conjecture of Turán from the 1940s asserts that the number of edges in an $n$-vertex 33-graph without the tetrahedron, the complete $3$-graph on four vertices, cannot exceed that of the balanced complete cyclic $3$-partite 33-graph, whose edges are of types $V_{1}V_{2}V_{3}$, $V_{1}V_{1}V_{2}$, $V_{2}V_{2}V_{3}$, and $V_{3}V_{3}V_{1}$. A recent surprising result of Balogh–Clemen–Lidický [J. Lond. Math. Soc. (2) 106 (2022)] shows that this conjecture is asymptotically true in the $\ell_{2}$-norm, where the number of edges is replaced by the sum of squared codegrees. They further conjectured that, in this $\ell_{2}$-norm setting, the $3$-partite construction is uniquely extremal for large $n$. We confirm this conjecture.
Two key ingredients in our proofs include establishing a Mantel theorem for vertex-colored graphs that forbid certain types of triangles, and introducing a novel procedure integrated into Simonovits’ stability method, which essentially reduces the task to verifying that the $\ell_{2}$-norm of certain near-extremal constructions increases under suitable local modifications. The strategy in the latter may be of independent interest and potentially applicable to other extremal problems.
footnotetext: Keywords: Tetrahedron Conjecture, hypergraph Turán problem, stability method
MSC2020: 05C35, 05C65, 05D05
1 Introduction
Given an integer $r\geq 2$, an $r$-uniform hypergraph (henceforth an $r$-graph) $\mathcal{H}$ is a collection of $r$-subsets of some set $V$. We call $V$ the vertex set of $\mathcal{H}$ and denote it by $V(\mathcal{H})$. When $V$ is understood, we usually identify a hypergraph $\mathcal{H}$ with its set of edges.
Given a family $\mathcal{F}$ of $r$-graphs, an $r$-graph $\mathcal{H}$ is $\mathcal{F}$-free if it does not contain any member of $\mathcal{F}$ as a subgraph. A fundamental problem in Extremal Combinatorics is to determine the extremal properties of $\mathcal{F}$-free $r$-graphs on $n$ vertices. In particular, the Turán problem asks for the Turán number $\mathrm{ex}(n,\mathcal{F})$, the maximum number of edges in an $\mathcal{F}$-free $r$-graphs on $n$ vertices.
For $r=2$ (i.e., graphs), the value of $\mathrm{ex}(n,\mathcal{F})$ is well understood up to an additive error term of order $o(n^{2})$ thanks to the celebrated Erdős–Stone Theorem [ES46] (see also [ES66]). In contrast, for $r\geq 3$, determining $\mathrm{ex}(n,\mathcal{F})$ even asymptotically is notoriously difficult in general. Denote by $K_{\ell}^{r}$ the complete $r$-graph on $\ell$ vertices (with the superscript omitted when $r=2$). The seminal work of Turán [Tur41] determined the exact value of $\mathrm{ex}(n,K_{\ell})$ for every $\ell\geq 3$ (with the case $\ell=3$ solved earlier by Mantel [Man07]). Erdős [Erd81] offered \$500 for the asymptotic determination of $\mathrm{ex}(n,K_{\ell}^{r})$ for any single pair $\ell,r$ with $\ell>r\geq 3$. This prize remains unclaimed, despite decades of intensive work. For further background and related results, we refer the reader to the surveys [Für91, dC94, Sid95, Kee11].
Perhaps the most famous instance is the tetrahedron $K_{4}^{3}$, the complete $3$-graph on four vertices. Given three pairwise disjoint sets $V_{1},V_{2},V_{3}$, let $\mathbb{C}[V_{1},V_{2},V_{3}]$ denote the $3$-graph on $V_{1}\cup V_{2}\cup V_{3}$ consisting of all triples of types $V_{1}V_{2}V_{3}$, $V_{1}V_{1}V_{2}$, $V_{2}V_{2}V_{3}$, and $V_{3}V_{3}V_{1}$ (see Figure 1), that is,
$\mathbb{C}[V_{1},V_{2},V_{3}]\coloneqq\left\{e\colon\big(|e\cap V_{1}|,|e\cap V_{2}|,|e\cap V_{3}|\big)\in\{(1,1,1),(2,1,0),(0,2,1),(1,0,2)\}\right\}.$
When $V_{1}\cup V_{2}\cup V_{3}=[n]$ is a balanced partition of $[n]$, we write $\mathbb{C}_{n}$ for $\mathbb{C}[V_{1},V_{2},V_{3}]$.
The famous Tetrahedron Conjecture of Turán [Tur41] asserts that $\mathrm{ex}(n,K_{4}^{3})=|\mathbb{C}_{n}|$. Various distinct constructions achieving the same bound as $\mathbb{C}_{n}$ were discovered later (see, e.g., [Bro83, Kos82, FDF88, Fro08]). In particular, this implies that the Turán problem for $K_{4}^{3}$ does not exhibit the stability property (see [LM22, Proposition 1.7]), which is likely one reason for the notorious difficulty of the conjecture. Successively better upper bounds for $\mathrm{ex}(n,K_{4}^{3})$ have been obtained by many researchers (see e.g., [dC88, CL99, Raz10a, BT11]), with the current record established by the computer-assisted flag algebra framework of Razborov [Raz07].
$V_{1}$$V_{2}$$V_{3}$$V_{1}$$V_{2}$
Figure 1: Structures of $\mathbb{C}[V_{1},V_{2},V_{3}]$ and $\mathbb{B}[V_{1},V_{2}]$.
Given an $r$-graph $\mathcal{H}$ on a vertex set $V$, the codegree of an (r1)(r-1)-subset $S\subseteq V$ in $\mathcal{H}$, denoted by $d_{\mathcal{H}}(S)$, is the number of edges in $\mathcal{H}$ containing $S$. An simple double-counting argument shows that $\sum_{e\in\binom{V}{r-1}}d_{\mathcal{H}}(e)=r|\mathcal{H}|$. Balogh–Clemen–Lidický viewed the summation on the left side of the equality as the $\ell_{1}$-norm of $\mathcal{H}$, and they further introduced the $\ell_{2}$-norm of $\mathcal{H}$, given by
$\lVert\mathcal{H}\rVert_{2}\coloneqq\sum_{e\in\binom{V}{r-1}}d_{\mathcal{H}}^{2}(e).$
The Turán number of $\mathcal{F}$ in the $\ell_{2}$-norm is then defined as
$\mathrm{ex}_{\ell_{2}}(n,\mathcal{F})\coloneqq\max\left\{\lVert\mathcal{H}\rVert_{2}\colon\text{$v(\mathcal{H})=n$ and $\mathcal{H}$ is $\mathcal{F}$-free}\right\}.$
The Turán density of $\mathcal{F}$ in the $\ell_{2}$-norm is given by
$\pi_{\ell_{2}}(\mathcal{F})\coloneqq\lim_{n\to\infty}\frac{\mathrm{ex}_{\ell_{2}}(n,\mathcal{F})}{\binom{n}{r-1}(n-r+1)^{2}}.$
The construction $\mathbb{C}_{n}$ shows that $\pi_{\ell_{2}}(K_{4}^{3})\geq\frac{1}{3}$. Using Razborov’s flag algebra framework, Balogh–Clemen–Lidický [BCL22] proved, with computer assistance, that $\pi_{\ell_{2}}(K_{4}^{3})\leq\frac{1}{3}$ and thus established the following theorem.
Theorem 1.1 ([BCL22, Theorem 1.2]).
We have $\mathrm{ex}_{\ell_{2}}(n,K_{4}^{3})=\frac{n^{4}}{6}+o(n^{4})$.
They further conjectured (see [BCL22, Conjecture 1.3]) that for sufficiently large $n$, the construction $\mathbb{C}_{n}$ is the unique extremal configuration for the Turán problem of $K_{4}^{3}$ in the $\ell_{2}$-norm, in contrast to the $\ell_{1}$-norm setting, where exponentially many extremal constructions may exist. Our main result confirms this conjecture.
Theorem 1.2.
There exists $n_{0}$ such that the following holds for every $n\geq n_{0}$. Suppose that $\mathcal{H}$ is an $n$-vertex $K_{4}^{3}$-free $3$-graph. Then $\lVert\mathcal{H}\rVert_{2}\leq\lVert\mathbb{C}_{n}\rVert_{2}$, and equality holds iff $\mathcal{H}\cong\mathbb{C}_{n}$.
In Section 2, we prove Turán-type results on vertex-colored graphs without certain types of triangles, in both the $\ell_{1}$-norm (Theorem 2.1) and the $\ell_{2}$-norm (Theorem 2.3), together with a corresponding stability theorem (Theorem 2.2). Although the $\ell_{2}$-norm result can in principle be proved without computer-assisted flag algebra computations, the argument we found requires extremely lengthy case analyses; therefore, we include a flag algebra proof here. In Section 3, we present the proof of Theorem 1.2. The proof uses Simonovits’ stability method [Sim68], combined with a novel procedure that reduces the problem to the much simpler task of verifying that the $\ell_{2}$-norm of certain near-extremal constructions increases under specific local changes (see Lemmas 3.6 and 3.7). We believe that this framework applies to other extremal problems as well. In Section 4, we include some concluding remarks.
2 Mantel’s theorem in vertex-colored graphs
In this section, we consider a variation of Mantel’s theorem in vertex-colored graphs.
Let $G$ be a graph on vertex set $V$, and let $V\coloneqq V_{1}\cup V_{2}\cup V_{3}$ be a partition of $V$ (equivalently, a $3$-coloring of the vertices of $G$). For a triple of vertices $\{u,v,w\}\subseteq V$, we say that $G[\{u,v,w\}]$ is a triangle of type $V_{i}V_{j}V_{k}$ for some $(i,j,k)\in[3]^{3}$ if $\{u,v,w\}$ induces a copy of $K_{3}$ in $G$ and $(u,v,w)\in V_{i}\times V_{j}\times V_{k}$. Denote by $\rho_{3}(G)$ the density of triangles of types $V_{1}V_{2}V_{3}$, $V_{1}V_{1}V_{2}$, $V_{2}V_{2}V_{3}$, and $V_{3}V_{3}V_{1}$ in $G$, that is, the probability that a uniformly chosen $3$-subset of $V$ induces a triangle of one of these types. We say that $G$ is cyclically triangle-free if $\rho_{3}(G)=0$.
For a vertex set $S\subseteq V$, we use $G[S]$ to denote the induced subgraph of $G$ on $S$. For pairwise disjoint sets $S_{1},\ldots,S_{k}\subseteq V$, we use $G[S_{1},\ldots,S_{k}]$ to denote the induced $k$-partite subgraph of $G$ with parts $S_{1},\ldots,S_{k}$.
Let $\Lambda[V_{1},V_{2},V_{3}]$ denote the graph on vertex set $V$ whose edge set is
$\left\{e\in\tbinom{V}{2}\colon(|e\cap V_{1}|,|e\cap V_{2}|,|e\cap V_{3}|)\in\{(1,1,0),(0,1,1),(0,0,2)\}\right\}.$
Observe that $\Lambda[V_{1},V_{2},V_{3}]$ is cyclically triangle-free. Suppose that $|V_{1}|=|V_{2}|=|V_{3}|=n$. Then straightforward calculations show that
$|\Lambda[V_{1},V_{2},V_{3}]|$ $=2n^{2}+\binom{n}{2}=\frac{5}{2}n^{2}-\frac{n}{2},$
$\lVert\Lambda[V_{1},V_{2},V_{3}]\rVert_{2}$ $=n\cdot n^{2}+n\cdot(2n)^{2}+n\cdot(2n-1)^{2}=9n^{3}-4n^{2}+n.$
Throughout this section, we assume that $G$ is a graph with vertex set $V$, and that $V=V_{1}\cup V_{2}\cup V_{3}$ is a fixed partition. We further assume that all indices are taken modulo $3$.
The main results of this section are as follows.
Theorem 2.1.
[…]
$|G|\leq\frac{5}{2}n^{2}+\varepsilon n^{2}.$
We will also use the following stability theorem.
Theorem 2.2.
[…]
Then, there exists $i\in[3]$ such that
$|G[V_{i}]|+|G[V_{i+1}]|+|G[V_{i},V_{i+2}]|\leq\varepsilon n^{2}.$
We say $G$ is locally maximal if there exists $i\in[3]$ such that
  1. (i)
    $|G[V_{i},V_{i+1}]|+|G[V_{i+2}]|\geq|G[V_{i},V_{i+2}]|+|G[V_{i}]|$, and
  2. (ii)
    $|G[V_{i+1},V_{i+2}]|+|G[V_{i+2}]|\geq|G[V_{i},V_{i+2}]|+|G[V_{i+1}]|$.
Theorem 2.3.
[…]
$\lVert G\rVert_{2}\leq 9n^{3}+\varepsilon n^{3}.$
We remark that the locally maximal assumption is needed only for the flag algebra proof, the alternative proof, which avoids flag algebras, does not require this assumption.
2.1 Proof of Theorem 2.1
[…]
Theorem 2.4.
Suppose that $|V_{i}|=n$ for $i\in[3]$ and $G$ is cyclically triangle-free. Then $|G|\leq 5n^{2}/2+5n$.
For convenience, we assume that $|V_{i}|=n$ for $i\in[3]$ throughout this subsection.
We use a slightly modified version of the Zykov symmetrization [Zyk49] for the proof of Theorem 2.4. Given distinct vertices $u,v\subseteq V$, denote by $G_{u\to v}$ the graph obtained from $G$ by symmetrizing $u$ to $v$, that is,
$G_{u\to v}\coloneqq\big(G\setminus\left\{uw\colon w\in N_{G}(u)\right\}\big)\cup\left\{uw\colon w\in N_{G}(v)\right\}.$
We say that $u$ and $v$ are equivalent, denoted $u\sim v$, if $u$ and $v$ lie in the same part $V_{i}$ for some $i\in[3]$ and $N_{G}(u)=N_{G}(v)$ (and hence $\{u,v\}\notin G$). For $i\in[3]$ and $v\in V_{i}$, write
$[v]\coloneqq\{u\in V_{i}\colon u\sim v\}$
for the equivalence class of $v$ in $V_{i}$ (with respect to $\sim$ in $G$).
We say that $G$ is locally symmetrized if, for each $i\in[3]$, any two nonadjacent vertices in $V_{i}$ are equivalent.
Observe that the symmetrization operation extends naturally to equivalence classes, that is, we may symmetrize all vertices in $[u]$ to a vertex $[v]$ simultaneously, denoted by $G_{[u]\to[v]}$. The advantage of this extension is that each application reduces the number of equivalence classes by one, thus, after finitely many steps, we obtain a locally symmetrized graph. Moreover, it is easy to verify that if $G$ is cyclically triangle-free, then $G_{[u]\to[v]}$ is also cyclically triangle-free. In addition, if follows from the definition that for every pair of equivalence classes in $V_{i}$ for some $i\in[3]$,
$|G|\leq\max\left\{|G_{[u]\to[v]}|,~|G_{[v]\to[u]}|\right\}.$
Therefore, in the proof of Theorem 2.4, we may assume that $G$ is locally symmetrized.
We associate to $G$ the mixed graph $\vec{G}$, obtained by orienting every edge between $V_{i}$ and $V_{i+1}$ from $V_{i}$ to $V_{i+1}$, while leaving edges inside each $V_{i}$ undirected.
A directed walk, path, or cycle in $\vec{G}$ is called distinct if no two of its vertices belong to the same equivalence class in $G$.
For $i\in[3]$ and $v\in V_{i}$, define its in-neighborhood, out-neighborhood, and internal neighborhood by
$N^{-}(v)\coloneqq N_{G}(v)\cap V_{i-1},\quad N^{+}(v)\coloneqq N_{G}(v)\cap V_{i+1},\quad N^{\mathrm{int}}(v)\coloneqq N_{G}(v)\cap V_{i}.$
The following fact follows directly from the definition of locally symmetrized.
Fact 2.5.
Suppose that $G$ is locally symmetrized. Then the following statements hold.
  1. (i)
    Each equivalence class is an independent set in $G$.
  2. (ii)
    For each $i\in[3]$ and every pair of distinct equivalence classes $x],[y]\subseteq V_{i}$, the induced bipartite subgraph $G\big[[x],[y]\big]$ is complete. In particular, for every vertex $v\in V_{i}$, we have $|N^{\mathrm{int}}(v)|=|V_{i}|-|[v]|$.
  3. (iii)
    Every minimum-length directed cycle in $\vec{G}$ is distinct.
The following fact follows directly from the definition of cyclically triangle-free and Fact 2.5.
Fact 2.6.
Suppose that $G$ is locally symmetrized and cyclically triangle-free. Then the following statements hold.
  1. (i)
    For every $i\in[3]$ and $v\in V_{i}$, either $N^{-}(v)=\emptyset$ or $N^{-}(v)=[u]$ for some equivalence class $u]\subseteq V_{i-1}$.
  2. (ii)
    Let $i\in[3]$. Suppose that $(u,v)\in V_{i}\times V_{i+1}$ is a directed edge in $\vec{G}$. Then $N^{-}(u)\cap N^{+}(v)=\emptyset$. In particular, $|N^{-}(u)\cup N^{+}(v)|\leq|V_{i-1}|$.
The following lemma is crucial for the proof of Theorem 2.4.
Lemma 2.7.
Suppose that $G$ is locally symmetrized and cyclically triangle-free. Let $W=x_{1}y_{1}z_{1}\ldots x_{k}y_{k}z_{k}$ be a distinct directed path in $\vec{G}$ with $(x_{i},y_{i},z_{i})\in V_{1}\times V_{2}\times V_{3}$ for $i\in[k]$. Suppose that either $N^{-}_{G}(x_{1})=\emptyset$ or $N^{-}_{G}(x_{1})=[z_{k}]$. Then
$\sum_{i\in[k]}\big(d_{G}(x_{i})+d_{G}(y_{i})+d_{G}(z_{i})\big)\leq 3(k+1)n.$
Proof of Lemma 2.7.
For $i\in[2,k]$, since $(z_{i-1},x_{i})\in\vec{G}$, it follows from Fact 2.6 (i) that $N^{-}(x_{i})=[z_{i-1}]$. Combining this with the assumption that $N^{-}_{G}(x_{1})=\emptyset$ or $N^{-}_{G}(x_{1})=[z_{k}]$, we obtain
$\sum_{i\in[k]}|N^{-}(x_{i})|\leq\sum_{i\in[k]}|[z_{i}]|.$ (1)
For $i\in[k]$, it follows from Fact 2.5 (ii) that $N^{\mathrm{int}}(x_{i})=|V_{1}|-|[x_{i}]|=n-|[x_{i}]|$. Therefore,
$\sum_{i\in[k]}|N^{\mathrm{int}}(x_{i})|=kn-\sum_{i\in[k]}|[x_{i}]|.$ (2)
It follows from Fact 2.6 (i) that each vertex $u\in V_{2}$ is adjacent to at most one equivalence class in $V_{1}$. Thus,
$\sum_{i\in[k]}|N^{+}(x_{i})|\leq|V_{2}|=n.$
Combining this with (1) and (2), we obtain
$\sum_{i\in[k]}d_{G}(x_{i})$ $=\sum_{i\in[k]}\left(|N^{+}(x_{i})|+|N^{-}(x_{i})|+|N^{\mathrm{int}}(x_{i})|\right)$
$\leq(k+1)n+\sum_{i\in[k]}|[z_{i}]|-\sum_{i\in[k]}|[x_{i}]|.$
Applying the same argument to $y_{1},\ldots,y_{k}$ and $z_{1},\ldots,z_{k}$ yields
$\sum_{i\in[k]}d_{G}(y_{i})$ $\leq(k+1)n+\sum_{i\in[k]}|[x_{i}]|-\sum_{i\in[k]}|[y_{i}]|,$
$\sum_{i\in[k]}d_{G}(z_{i})$ $\leq(k+1)n+\sum_{i\in[k]}|[y_{i}]|-\sum_{i\in[k]}|[z_{i}]|.$
Summing these three inequalities gives
$\sum_{i\in[k]}\big(d_{G}(x_{i})+d_{G}(y_{i})+d_{G}(z_{i})\big)$ $\leq 3(k+1)n,$
which proves Lemma 2.7.
We are now ready to present the proof of Theorem 2.4.
Proof of Theorem 2.4.
Suppose that $G$ is cyclically triangle-free and has the maximum possible number of edges. By the discussion above, we may assume that $G$ is locally symmetrized. Suppose to the contrary that $|G|>5n^{2}/2+5n$.
Claim 2.8.
Every triple of vertices $(x,y,z)\in V_{1}\times V_{2}\times V_{3}$ satisfies
$d_{G}(x)+d_{G}(y)+d_{G}(z)>5n.$ (3)
Proof of Claim 2.8.
Suppose to the contrary that $d_{G}(x)+d_{G}(y)+d_{G}(z)\leq 5n$. Let us label vertices in $V_{1},V_{2},V_{3}$ by $V_{1}=\{x_{1},\ldots,x_{n}\}$, $V_{2}=\{y_{1},\ldots,y_{n}\}$, and $V_{3}=\{z_{1},\ldots,z_{n}\}$ with $(x,y,z)=(x_{1},y_{1},z_{1})$. Since
$\sum_{i\in[n]}\big(d_{G}(x_{i})+d_{G}(y_{i})+d_{G}(z_{i})\big)=2|G|\geq 5n^{2}+10n,$
by averaging, there exists a triple $(x_{i_{0}},y_{i_{0}},z_{i_{0}})$ such that
$d_{G}(x_{i_{0}})+d_{G}(y_{i_{0}})+d_{G}(z_{i_{0}})\geq 5n+10.$
Let $\tilde{G}$ be the graph obtained from $G$ by first removing all edges incident to $\{x_{1},y_{1},z_{1}\}$, and then symmetrizing $x_{1}$ to $x_{i_{0}}$, $y_{1}$ to $y_{i_{0}}$, and $z_{1}$ to $z_{1_{0}}$. Note that $\tilde{G}$ remains cyclically triangle-free. However,
[…]
$\geq|G|-5n+(5n+10-9)>|G|,$
contradicting the maximality of $G$.
Claim 2.9.
The mixed graph $\vec{G}$ does not contain any directed cycle. Consequently, every directed walk in $\vec{G}$ is a directed path.
Proof of Claim 2.9.
Suppose to the contrary that this claim fails. Let $C=x_{1}y_{1}z_{1}\cdots x_{k}y_{k}z_{k}$ be a shortest directed cycle in $\vec{G}$. By relabeling the vertices of $C$ if necessary, we may assume that $(x_{i},y_{i},z_{i})\in V_{1}\times V_{2}\times V_{3}$ for $i\in[k]$. Since $G$ does not contain triangles of type $V_{1}V_{2}V_{3}$, we have $k\geq 2$. Moreover, because $C$ s chosen to be shortest, it must be distinct (otherwise, we obtain a shorter directed cycle).
Applying Lemma 2.7 to the path $x_{1}y_{1}z_{1}\cdots x_{k}y_{k}z_{k}$, we obtain
$\sum_{i\in[k]}\big(d_{G}(x_{i})+d_{G}(y_{i})+d_{G}(z_{i})\big)\leq 3(k+1)n.$
By averaging, there exists an index $i_{\ast}\in[k]$ such that
$d_{G}(x_{i_{\ast}})+d_{G}(y_{i_{\ast}})+d_{G}(z_{i_{\ast}})\leq 3\left(1+\frac{1}{k}\right)n\leq 3\left(1+\frac{1}{2}\right)n<5n,$
a contradiction to (3).
Now fix a longest directed walk $P=w_{1}\cdots w_{m}$ of $\vec{G}$. It follows from Claim 2.9 that $P$ is a path.
Claim 2.10.
We have $m\geq 3$.
Proof of Claim 2.10.
Suppose to the contrary that $m\leq 2$. If $m=1$, then $G[V_{i},V_{j}]=\emptyset$ for every $\{i,j\}\in\tbinom{[3]}{2}$. In this case,
$|G|=|G[V_{1}]|+|G[V_{2}]|+|G[V_{3}]|\leq 3\binom{n}{2}<\frac{5n^{2}}{2}+5n,$
a contradiction. Thus, we may assume that $m=2$.
By symmetry, we may assume that $w_{1}\in V_{1}$ and $w_{2}\in V_{2}$. Fix an arbitrary vertex $w_{3}\in V_{3}$. It follows from the maximality of $P$ that
$N^{-}(w_{1})=N^{+}(w_{2})=\emptyset\quad\text{and}\quad N^{+}(w_{1})\cap N^{-}(w_{3})=\emptyset.$
Combining it with Fact 2.5 (ii) and Fact 2.6 (i), we obtain
$\sum_{i\in[3]}d_{G}(w_{i})$ $=\sum_{i\in[3]}\left(|N^{-}(w_{i})|+|N^{\mathrm{int}}(w_{i})|+|N^{+}(w_{i})|\right)$
$\leq\big(n-|[w_{1}]|+|N^{+}(w_{1})|\big)+\big(|[w_{1}]|+n-|[w_{2}]|\big)+\big(|N^{-}(w_{3})|+2n\big)$
$\leq 4n+|N^{+}(w_{1})|+|N^{-}(w_{3})|-|[w_{2}]|\leq 5n-|[w_{2}]|<5n,$
a contradiction to (3).
We now consider the last three vertices $\{w_{m-2},w_{m-1},w_{m}\}$ of the path $P$. By symmetry, we may assume that $(w_{m-2},w_{m-1},w_{m})\in V_{1}\times V_{2}\times V_{3}$. For convenience, set $(x,y,z)\coloneqq(w_{m-2},w_{m-1},w_{m})$. Since $P$ is a longest directed path, we have $N^{+}(z)=\emptyset$. Moreover, as $(x,y)$ is a directed edge, it follows from Fact 2.6 (ii) that $|N^{-}(x)|+|N^{+}(y)|\leq|V_{3}|=n$. Therefore,
$d_{G}(x)+d_{G}(y)+d_{G}(z)$
$\qquad\leq\big(|N^{-}(x)|+n-|[x]|+n\big)+\big(|[x]|+n-|[y]|+|N^{+}(y)|\big)+\big(|[y]|+n\big)$
$\qquad=4n+|N^{-}(x)|+|N^{+}(y)|\leq 5n,$
a contradiction to (3). This completes the proof of Theorem 2.1.
2.2 Proof of Theorem 2.2
In this subsection, we present the proof of Theorem 2.2. We will use the classical result of Bollobás–Erdős–Straus [BES74], together with a stability version recently established in [CLY22]. For our purposes, we state slightly modified variants of these results, which follow straightforwardly from the Removal Lemma.
Theorem 2.11 ([BES74] and [CLY22, Theorem 1.5]).
[…]
$|H|\leq 2n^{2}+\xi n^{2}.$
[…]
$|H\triangle K[U_{i1}\cup U_{i+1},U_{i2}\cup U_{i+2}]|\leq\xi n^{2}.$
Proof of Theorem 2.2.
Fix $\varepsilon>0$. Let $\delta,\delta_{1},\delta_{2}>0$ be sufficiently small such that $\delta\ll\delta_{1}\ll\delta_{2}\ll\varepsilon$. Let $n$ be sufficiently large. Let $V_{1},V_{2},V_{3}$ be pairwise disjoin sets satisfying $\max_{i\in[3]}\left||V_{i}|-n\right|\leq\delta n$. Suppose that $G$ satisfies $\rho_{3}(G)\leq\delta$ and
$|G|+|G[V_{1},V_{2},V_{3}]|\geq\frac{9n^{2}}{2}-\delta n^{2}.$ (4)
We can choose $\delta$ sufficiently small such that, by Theorems 2.1 and 2.11,
$\max\left\{|G|-\frac{5n^{2}}{2},~|G[V_{1},V_{2},V_{3}]|-2n^{2}\right\}\leq\delta_{1}n^{2},$ (5)
and moreover, such that, by the Removal Lemma, there exists a subgraph $\tilde{G}\subseteq G$ with $\rho_{3}(\tilde{G})=0$ and $|\tilde{G}|\geq|G|-\delta_{1}n^{2}$. Combining this with (4) and (5), we obtain
$\min\left\{|\tilde{G}|-\frac{5n^{2}}{2},~|\tilde{G}[V_{1},V_{2},V_{3}]|-2n^{2}\right\}\geq-3\delta_{1}n^{2}.$ (6)
Applying Theorem 2.11 to the $3$-partite graph $\tilde{G}[V_{1},V_{2},V_{3}]$, we conclude that there exist $i\in[3]$ and a partition $V_{i1}\cup V_{i2}=V_{i}$ such that
$|\tilde{G}[V_{1},V_{2},V_{3}]\triangle K[V_{i1}\cup V_{i+1},V_{i2}\cup V_{i+2}]|\leq\delta_{2}n^{2}.$ (7)
By symmetry, we may assume that $i=1$.
Claim 2.12.
The following statements hold.
  1. (i)
    $\max\left\{|\tilde{G}[V_{11},V_{2}]|,~|\tilde{G}[V_{12},V_{3}]|,~|\tilde{G}[V_{2}]|,~|\tilde{G}[V_{12}]|\right\}\leq\delta_{2}n^{2}$.
  2. (ii)
    $|V_{11}|\leq\delta_{2}^{1/2}n$ or $|\tilde{G}[V_{3}]|\leq\delta_{2}^{1/2}n^{2}$.
Proof of Claim 2.12.
Let $M\coloneqq K[V_{11}\cup V_{2},V_{12}\cup V_{3}]\setminus\tilde{G}[V_{1},V_{2},V_{3}]$. It follows from (7) that $|M|\leq\delta_{2}n^{2}$. First, we prove that $|\tilde{G}[V_{11},V_{2}]|\leq\delta_{2}n^{2}$. Let
$\mathcal{T}\coloneqq\big\{\{u,v,w\}\colon\text{$uv\in\tilde{G}[V_{11},V_{2}]$, $w\in V_{3}$, and $\{uw,vw\}\cap M\neq\emptyset$}\big\}.$
Fix $uv\in\tilde{G}[V_{11},V_{2}]$. Since $\tilde{G}$ does not contain triangles of type $V_{1}V_{2}V_{3}$, for every $w\in V_{3}$ we have $\{uw,vw\}\cap M\neq\emptyset$. It follows that $|\mathcal{T}|\geq|\tilde{G}[V_{11},V_{2}]||V_{3}|$. On the other hand, every pair $uw\in M$ is contained in at most $n$ members of $\mathcal{T}$. Thus, we have $|\mathcal{T}|\leq n|M|$. Combining the two bounds, we obtain
$|\tilde{G}[V_{11},V_{2}]|\leq\frac{n|M|}{|V_{3}|}=|M|\leq\delta_{2}n^{2}.$
The proofs for the bounds on $|\tilde{G}[V_{12},V_{3}]|$, $|\tilde{G}[V_{2}]|$ (here we use the assumption that $\tilde{G}$ contains no triangles of type $V_{2}V_{2}V_{3}$), $|\tilde{G}[V_{12}]|$ (here we use the assumption that $\tilde{G}$ contains no triangles of type $V_{1}V_{1}V_{2}$), and for Claim 2.12 (ii) (here we use the assumption that $\tilde{G}$ contains no triangles of type $V_{3}V_{3}V_{1}$) are similar, thus we omit them here.
Suppose that $|V_{11}|\leq\delta_{2}^{1/2}n$. Then it follows from Claim 2.12 (i) that
$|\tilde{G}[V_{1}]|+|\tilde{G}[V_{2}]|+|\tilde{G}[V_{1},V_{3}]|$
$\qquad=|\tilde{G}[V_{11}]|+|\tilde{G}[V_{11},V_{12}]|+|\tilde{G}[V_{12}]|+|\tilde{G}[V_{2}]|+|\tilde{G}[V_{11},V_{3}]|+|\tilde{G}[V_{12},V_{3}]|$
$\qquad\leq\tbinom{|V_{11}|}{2}+|V_{11}|n+\delta_{2}n^{2}+\delta_{2}n^{2}+|V_{11}|n+\delta_{2}n^{2}\leq 4\delta_{2}n^{2}+2\delta_{2}^{1/2}n^{2}\leq\varepsilon n^{2},$
as desired.
Suppose that $|V_{11}|>\delta_{2}^{1/2}n$. Then it follows from Claim 2.12 (ii) that $|\tilde{G}[V_{3}]|\leq\delta_{2}^{1/2}n^{2}$. Combining this with Claim 2.12 (i), we obtain
$|\tilde{G}[V_{1}]|+|\tilde{G}[V_{2}]|+|\tilde{G}[V_{3}]|$
$\qquad=|\tilde{G}[V_{11}]|+|\tilde{G}[V_{11},V_{12}]|+\tilde{G}[V_{12}]+|\tilde{G}[V_{2}]|+|\tilde{G}[V_{3}]|$
$\qquad\leq\tbinom{|V_{1}|}{2}-\tbinom{|V_{12}|}{2}+\delta_{2}n^{2}+\delta_{2}n^{2}+\delta_{2}n^{2}\leq\frac{n^{2}}{2}+3\delta_{2}^{1/2}n^{2}-\frac{|V_{12}|^{2}}{2}.$ (8)
On the other hand, it follows from (5) and (6)
$|\tilde{G}[V_{1}]|+|\tilde{G}[V_{2}]|+|\tilde{G}[V_{3}]|$ $=|\tilde{G}|-|\tilde{G}[V_{1},V_{2},V_{3}]|$
$\geq\frac{5}{2}n^{2}-3\delta_{1}n^{2}-\left(2n^{2}+\delta_{1}n^{2})\right)\geq\frac{1}{2}n^{2}-4\delta_{1}n^{2}.$
Combining this with (2.2), we obtain
$\frac{|V_{12}|^{2}}{2}\leq 3\delta_{2}^{1/2}n^{2}+4\delta_{1}n^{2}\leq 4\delta_{2}^{1/2}n^{2},$
which implies that $|V_{12}|\leq 2\sqrt{2}\delta_{2}^{1/4}n$. Consequently,
$|\tilde{G}[V_{2}]|+|\tilde{G}[V_{3}]|+|\tilde{G}[V_{1},V_{2}]|$
$\qquad=|\tilde{G}[V_{2}]|+|\tilde{G}[V_{3}]|+|\tilde{G}[V_{11},V_{2}]|+|\tilde{G}[V_{12},V_{2}]|$
$\qquad=\delta_{2}n^{2}+\delta_{2}^{1/2}n^{2}+\delta_{2}n^{2}+2\sqrt{2}\delta_{2}^{1/4}n^{2}\leq\varepsilon n^{2},$
as desired. This completes the proof of Theorem 2.2.
2.3 Proof of Theorem 2.3
In this section, we present the proof of Theorem 2.3, derived by computer using the flag algebra method of Razborov [Raz07], which is also described in e.g. [Raz10b, BaberTalbot11, dCSdOFS16, GGH+22]. Since this method is well-known by now, we will be very brief. In particular, we omit many definitions, referring the reader to [Raz07, Raz10b] for any missing notions. Roughly speaking, a flag algebra proof using 0-flags on $m$ vertices of an upper bound $u\in\mathbb{R}$ on the given objective function $f$ consists of an identity
$u-f(\mathcal{H})=\mathrm{SOS}+\sum_{F\in\mathcal{F}_{m}^{0}}c_{F}\cdot p(F,\mathcal{H})+o(1),$
which is asymptotically true for any admissible $\mathcal{H}$ with $|V(\mathcal{H})|\to\infty$, where the $\mathrm{SOS}$-term can be represented as a sum of squares (as described e.g. in [Raz10b, Section 3]), each coefficient $c_{F}\in\mathbb{R}$ is non-negative, and $\mathcal{F}_{m}^{0}$ consists of isomorphism types of 0-flags (unlabeled $3$-graphs) with $m$ vertices. If $f(\mathcal{H})$ can be represented as a linear combination of the densities of members of $\mathcal{F}_{m}^{0}$ in $\mathcal{H}$ then finding the smallest possible $u$ amounts to solving a semi-definite program (SDP) with $|\mathcal{F}_{m}^{0}|$ linear constraints (so we write the size of $\mathcal{F}_{m}^{0}$ in each case to give the reader some idea of the size of the programs that we had to solve).
We formed the corresponding SDPs and then analyzed the solutions returned by computer, using a modified version of the SageMath package, which can be found in the GitHub repository https://github.com/bodnalev/sage. The script together with the certificate can be found at https://github.com/xliu2022/xliu2022.github.io/tree/main/FlagAlgebra_Certificate/ThreeColored-Mantel-L2norm.
Proof of Theorem 2.3.
Suppose to the contrary that this theorem fails. Then there exists a constant $\varepsilon>0$, a decreasing infinite sequence $\left(\delta_{i}\right)_{i=1}^{\infty}$ of positive real numbers with $\lim_{i\to\infty}\delta_{i}=0$, and an increasing (i.e., the number of vertices is strictly increasing) infinite sequence $\left(G_{i}\right)_{i=1}^{\infty}$ of graphs such that
  1. (i)
    $G_{i}$ is a graph on $V^{i}=V_{1}^{i}\cup V_{2}^{i}\cup V_{3}^{i}$, where $V_{1}^{i},V_{2}^{i},V_{3}^{i}$ are pairwise disjoint and satisfy
    $\left|V_{j}^{i}-\tfrac{|V^{i}|}{3}\right|\leq\delta_{i}|V^{i}|\quad\text{for}\quad j\in[3].$
  2. (ii)
    $G_{i}$ is locally maximal with respect to $(V_{1}^{i},V_{2}^{i},V_{3}^{i})$, that is, up to a cyclic permutation of the sub-indices,
    $|G_{i}[V_{1}^{i},V_{2}^{i}]|+|G_{i}[V_{3}^{i}]|$ $\geq|G_{i}[V_{1}^{i},V_{3}^{i}]|+|G_{i}[V_{1}^{i}]|,$
    $|G_{i}[V_{2}^{i},V_{3}^{i}]|+|G_{i}[V_{3}^{i}]|$ $\geq|G_{i}[V_{1}^{i},V_{3}^{i}]|+|G_{i}[V_{2}^{i}]|.$
  3. (iii)
    $\rho_{3}(G_{i})\leq\delta_{i}$.
  4. (iv)
    $\lVert G_{i}\rVert_{2}\geq(1/3+\varepsilon)|V^{i}|^{3}$.
We would like to run flag algebra calculations on the limit $\phi$ of the sequence $\left(G_{i}\right)_{i=1}^{\infty}$ in the theory of $3$-colored graphs, that is, we have three unary relations $U_{1},U_{2},U_{3}$ with the only restriction that each vertex in a flag satisfies exactly one of them, equivalently, these unitary relations encode a $3$-partition of the vertex set.
For $i\in[3]$, let $v_{i}$ denote the untyped single-vertex flag whose vertex has color $i$. Then, in the limit, part (i) translates to the statement that
$\phi(v_{i})=1/3\quad\text{for}\quad i\in[3].$ (9)
For $(i,j)\in[3]^{2}$, let $E_{i,j}$ denote the untyped two-vertex flag with one vertex colored $i$ and the other vertex colored $j$. Then, in the limit, part (ii) translates to the statement that
$\phi(E_{1,2})+\phi(E_{3,3})-\phi(E_{1,3})-\phi(E_{1,1})$ $\geq 0,$ (10)
$\phi(E_{2,3})+\phi(E_{3,3})-\phi(E_{1,3})-\phi(E_{2,3})$ $\geq 0.$
For $(i,j,k)\in[3]^{3}$, let $T_{i,j,k}$ denote the untyped three-vertex flag with one vertex colored $i$, one vertex colored $j$, and the remaining vertex colored $k$. Then, in the limit, part (iii) translates to the statement that
$\phi(T_{1,2,3})+\phi(T_{1,1,2})+\phi(T_{2,2,3})+\phi(T_{3,3,1})=0.$ (11)
Thus, we can restrict the flag algebra calculations to the theory of $\{T_{1,2,3},T_{1,1,2},T_{2,2,3},T_{3,3,1}\}$-free $3$-colored graphs.
Let $S_{2}$ denote the two-edge star and $K_{3}$ denote the triangle. For a graph $F$, let $\rho_{\mathrm{ind}}(F,G_{i})$ denote the density of induced copies of $F$ in $G_{i}$, i.e., the probability that the induced subgraph of $G$ on a uniformly sampled $v(F)$-subset of $V(G)$ is isomorphic to $F$. By definition,
$\lVert G_{i}\rVert_{2}=\sum_{v\in V^{i}}d_{G_{i}}^{2}(v)$ $=\sum_{v\in V^{i}}\left(2\tbinom{d_{G_{i}}(v)}{2}+d_{G_{i}}(v)\right)$
$=2\big(\rho_{\mathrm{ind}}(S_{2},G_{i})+3\rho_{\mathrm{ind}}(K_{3},G_{i})\big)\tbinom{|V^{i}|}{3}+2|G_{i}|.$
Let $\mathcal{S}$ denote the collection of untyped three-vertex flag with exactly two edges and arbitrary vertex colors (there are $18$ such flags). Let $\mathcal{K}$ denote the collection of untyped three-vertex flag with three edges and arbitrary vertex colors. Then, by the equality above, part (iv) translates, in the limit, to the statement that
$\sum_{S\in\mathcal{S}}\phi(S)+3\sum_{K\in\mathcal{K}}\phi(K)\geq 1+3\varepsilon.$ (12)
We can now run the usual flag algebra calculations in the theory of $\{T_{1,2,3},T_{1,1,2},T_{2,2,3},T_{3,3,1}\}$-free $3$-colored graphs, where each of the inequalities in (9) to (10) can be multiplied by an unknown non-negative combination of respectively 0-flags. The final inequality should prove (12).
Our computation uses 55-vertex flags (where $|\mathcal{F}_{5}^{0}|=1968$). The output from the computer shows that
$\sum_{S\in\mathcal{S}}\phi(S)+3\sum_{K\in\mathcal{K}}\phi(K)\leq 1,$
which contradicts (12). This completes the proof of Theorem 2.3.
3 Proof of Theorem 1.2
3.1 Preliminaries
Given pairwise disjoint sets $V_{1},\ldots,V_{t}$, denote by $K[V_{1},\ldots,V_{k}]$ the complete $k$-partite graph with parts $V_{1},\ldots,V_{t}$. Denote by $\overline{K}[V_{1},\ldots,V_{k}]$ the complement of $K[V_{1},\ldots,V_{k}]$, that is,
$\overline{K}[V_{1},\ldots,V_{k}]\coloneqq\tbinom{V_{1}}{2}\cup\cdots\cup\tbinom{V_{k}}{2}$.
Let $\mathcal{H}$ be a $3$-graph on vertex set $V$. The shadow of $\mathcal{H}$ is defined as
$\partial\mathcal{H}\coloneqq\left\{e\in\tbinom{V}{2}\colon\text{there exists $E\in\mathcal{H}$ such that $e\subseteq E$}\right\}.$
The link of a vertex $v\in V$ in $\mathcal{H}$ is given by
$L_{\mathcal{H}}(v)\coloneqq\big\{e\in\partial\mathcal{H}\colon\{v\}\cup e\in\mathcal{H}\big\}.$
The degree of $v$ is $d_{\mathcal{H}}(v)\coloneqq|L_{\mathcal{H}}(v)|$. We use $\delta(\mathcal{H})$, $d(\mathcal{H})$, and $\Delta(\mathcal{H})$ to denote the minimum, average, and maximum degree of $\mathcal{H}$, respectively.
The neighborhood of $v$ in $\mathcal{H}$ is defined as
$N_{\mathcal{H}}(v)\coloneqq\big\{u\in V\setminus\{v\}\colon\text{there exists $E\in\mathcal{H}$ such that $\{u,v\}\subseteq E$}\big\}.$
It follows from the definition that $u\in N_{\mathcal{H}}(v)$ if and only if $u\neq v$ and $d_{\mathcal{H}}(uv)\neq 0$. Moreover, for every $u\in V\setminus\{v\}$, the degree of $u$ in the link graph $L_{\mathcal{H}}(v)$ is the same as the codegree $d_{\mathcal{H}}(uv)$.
For a $2$-subset $e\subseteq V$, the neighborhood of $e$ in $\mathcal{H}$ is defined as
$N_{\mathcal{H}}(e)\coloneqq\big\{v\in V\colon e\cup\{v\}\in\mathcal{H}\big\}.$
Recall that the codegree of $e$ in $\mathcal{H}$ is the same as the size of $N_{\mathcal{H}}(e)$.
For a vertex subset $S\subseteq V(\mathcal{H})$, we use $\mathcal{H}[S]$ to denote the induced subgraph of $\mathcal{H}$ on $S$, and use $\mathcal{H}-S$ to denote the induced subgraph of $\mathcal{H}$ on $V(\mathcal{H})\setminus S$. If $S=\{v\}$ consists of a single vertex, then we write $\mathcal{H}-v$ instead of $\mathcal{H}-\{v\}$ for simplicity.
Following the definition in [CIL+24], the $2$-norm degree of a vertex $v\in V$ is defined as
$s_{\mathcal{H}}(v)\coloneqq\lVert\mathcal{H}\rVert_{2}-\lVert\mathcal{H}-v\rVert_{2}.$
Straightforward calculations (see [CIL+24, Lemma 3.1]) show that
$s_{\mathcal{H}}(v)$ $=\lVert L_{\mathcal{H}}(v)\rVert_{2}+2\sum_{e\in L_{\mathcal{H}}(v)}d_{\mathcal{H}}(e)-d_{\mathcal{H}}(v).$ (13)
Denote by $s(\mathcal{H})$ the average $2$-norm degree of $\mathcal{H}$.
For convenience, we will omit the subscript $\mathcal{H}$ from the above notations when it is clear from the context.
For the remainder of this section, all indices are taken modulo $3$.
The following statements can be verified by straightforward calculations.
Fact 3.1.
Let $n\geq 6$ be an integer, and let $V_{1}\cup V_{2}\cup V_{3}=[n]$ be a partition of $[n]$. Then
$\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\leq\lVert\mathbb{C}_{n}\rVert_{2}.$
Fact 3.2.
Let $V_{1},V_{2},V_{3}$ be three pairwise disjoint sets with $|V_{i}|=x_{i}n$ for $i\in[3]$. Then
$\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\\ =\sum_{i\in[3]}\left(\tfrac{x_{i}x_{i+1}\left(2\left(x_{i}+x_{i+2}\right)^{2}+x_{i}x_{i+1}\right)}{2}n^{4}-\tfrac{x_{i}x_{i+1}\left(4x_{i}+x_{i+1}+4x_{i+2}\right)}{2}n^{3}+x_{i}x_{i+1}n^{2}\right).$ (14)
In particular, let $\delta\in(0,1/3)$ and suppose $n\geq\frac{9}{2\delta^{2}}$. If $x_{i}\geq 1/3-\delta$ for $i\in[3]$. Then
$\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\geq\frac{n^{4}}{6}-2\delta n^{4}.$
Proof of Fact 3.2.
Equality (14) follows directly from the definition of $\mathbb{C}[V_{1},V_{2},V_{3}]$ and the definition of $2$-norm. So it suffices to prove the second part of the statement. Suppose that $x_{i}\geq 1/3-\delta$ for $i\in[3]$. Then
$\frac{x_{i}x_{i+1}\left(2\left(x_{i}+x_{i+2}\right)^{2}+x_{i}x_{i+1}\right)}{2}n^{4}$ $\geq\frac{n^{4}}{18}-\frac{2\delta n^{4}}{3}+\left(3\delta^{2}-6\delta^{3}+\frac{9\delta^{4}}{2}\right)n^{4}$
$\geq\frac{n^{4}}{18}-\frac{2\delta n^{4}}{3}+\delta^{2}n^{4},$
where the last inequality uses the fact that $2\delta^{2}-6\delta^{3}+\frac{9\delta^{4}}{2}\geq 0$ for $\delta\in[0,1/3]$.
On the other hand, plugging in $x_{1}=x_{2}=x_{3}=1$, we obtain the trivial upper bound
$\frac{x_{i}x_{i+1}\left(4x_{i}+x_{i+1}+4x_{i+2}\right)}{2}n^{3}\leq\frac{9n^{3}}{2}.$
Combining this with the inequality above, we obtain
$\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\geq 3\left(\frac{n^{4}}{18}-\frac{2\delta n^{4}}{3}+\delta^{2}n^{4}-\frac{9n^{3}}{2}\right)\geq\frac{n^{4}}{6}-2\delta n^{4},$
completing the proof of Fact 3.2.
We will use the following inequality, which can be established by standard calculus, for convenience, we verified it using Mathematica [^2] A Mathematica notebook containing the calculations is available at https://github.com/xliu2022/xliu2022.github.io/blob/main/Mathematica_Calculations/TetrahedronL2Norm.nb. .
Lemma 3.3.
Suppose that $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$ are nonnegative real numbers satisfying $x_{1}+x_{2}+x_{3}=1$. Then
$x_{1}x_{2}x_{3}+\frac{x_{1}^{2}x_{2}}{2}+\frac{x_{2}^{2}x_{3}}{2}+\frac{x_{3}^{2}x_{1}}{2}\leq\frac{5}{54}-\frac{1}{50}\left(\left(x_{1}-\tfrac{1}{3}\right)^{2}+\left(x_{2}-\tfrac{1}{3}\right)^{2}+\left(x_{3}-\tfrac{1}{3}\right)^{2}\right).$
Since the shadow of $K_{4}^{3}$ is a complete graph, duplicating a vertex in a $K_{4}^{3}$-free $3$-graph does not create any copy of $K_{4}^{3}$. Thus, a simple deletion–duplication argument (see e.g. [HLZ25, Lemma 2.6]) shows that every extremal (in terms of $\ell_{2}$-degree) $K_{4}^{3}$-free $3$-graph is early regular in the $\ell_{2}$-norm degree.
Lemma 3.4.
Suppose that $\mathcal{H}$ is an $K_{4}^{3}$-free $n$-vertex 33-graph with $\lVert\mathcal{H}\rVert_{2}=\mathrm{ex}_{\ell_{2}}(n,K_{4}^{3})$. Then for every pair of vertices $\{u,v\}\subseteq V(\mathcal{H})$, we have $\left|s_{\mathcal{H}}(u)-s_{\mathcal{H}}(v)\right|\leq 60n^{2}$. In particular, for every vertex $v\in V(\mathcal{H})$,
$\left|s_{\mathcal{H}}(v)-s(\mathcal{H})\right|\leq 60n^{2}.$
The following stability theorem of Balogh–Clemen–Lidický will be crucial for our proof.
Theorem 3.5 ([BCL22, Theorem 1.5]).
[…]
$|\mathcal{H}\triangle\mathbb{C}_{n}|\leq\varepsilon n^{3}.$
3.2 Two key lemmas
In this subsection, we establish two lemmas providing local modifications (see Figure 2) of $3$-graphs “close” to $\mathbb{C}_{n}$ that increase their $\ell_{2}$-norm. In these lemmas, we do not require the $K_{4}^{3}$-freeness.
Let $\mathcal{H}$ be a $3$-graph. For every pair $e\subseteq V(\mathcal{H})$, define
$\mathcal{H}(e)\coloneqq\left\{E\in\mathcal{H}\colon e\subseteq E\right\}.$
Given a partition $V_{1}\cup V_{2}\cup V_{3}=V(\mathcal{H})$, let
$\mathcal{B}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\coloneqq\mathcal{\mathcal{H}}\setminus\mathbb{C}[V_{1},V_{2},V_{3}]\quad\text{and}\quad\mathcal{M}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\coloneqq\mathbb{C}[V_{1},V_{2},V_{3}]\setminus\mathcal{H}.$ (15)
We refer to $\mathcal{B}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ and $\mathcal{M}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ as the sets of bad edges and missing edges of $\mathcal{H}$ (with respects to $(V_{1},V_{2},V_{3})$).
We further partition the set of bad edges by setting
$\mathcal{B}^{\mathrm{int}}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ $\coloneqq\left\{e\in\mathcal{B}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\colon\text{$e\subseteq V_{i}$ for some $i\in[3]$}\right\},$ (16)
$\mathcal{B}^{\mathrm{bi}}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ $\coloneqq\mathcal{B}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\setminus\mathcal{B}^{\mathrm{int}}_{\mathcal{H}}[V_{1},V_{2},V_{3}].$
Partition the set of missing edges by setting
$\mathcal{M}^{\mathrm{tri}}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ $\coloneqq\left\{e\in\mathcal{M}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\colon\text{$|e\cap V_{i}|=1$ for $i\in[3]$}\right\},$ (17)
$\mathcal{M}^{\mathrm{bi}}_{\mathcal{H}}[V_{1},V_{2},V_{3}]$ $\coloneqq\mathcal{M}_{\mathcal{H}}[V_{1},V_{2},V_{3}]\setminus\mathcal{M}^{\mathrm{tri}}_{\mathcal{H}}[V_{1},V_{2},V_{3}].$
We will omit the subscript $\mathcal{H}$ and the bracket $V_{1},V_{2},V_{3}]$ if it is clear from the context.
$V_{1}$$V_{2}$$V_{3}$$V_{1}$$V_{2}$$V_{3}$
Figure 2: Bad and missing edges in Lemmas 3.6 and 3.7.
Lemma 3.6.
[…]
  1. (i)
    $\max_{i\in[3]}|V_{i}-n/3|\leq\xi n$,
  2. (ii)
    $\max\left\{\Delta(\mathcal{M}),~\Delta(\mathcal{B})\right\}\leq\xi n^{2}$,
  3. (iii)
    $d_{\mathcal{M}}(e_{\ast})\geq 47\xi^{1/2}n$,
  4. (iv)
    $d_{\mathcal{M}}(e_{\ast})\geq d_{\mathcal{B}}(e_{\ast})-\xi n$,
  5. (v)
    $d_{\mathcal{B}^{\mathrm{bi}}}(e_{\ast})\leq\xi n$.
Then the new $3$-graph $\mathcal{G}^{\ast}\coloneqq\left(\mathcal{G}\setminus\mathcal{B}(e_{\ast})\right)\cup\mathcal{M}(e_{\ast})$ satisfies
$\lVert\mathcal{G}^{\ast}\rVert_{2}>\lVert\mathcal{G}\rVert_{2}.$
Proof of Lemma 3.6.
Let $\mathcal{G}$, $(V_{1},V_{2},V_{3})$, and $e_{\ast}$ be as in the statement of the lemma. By symmetry, we may assume that $e_{\ast}=\{u_{1},u_{2}\}\subseteq V_{1}$. Let
$\mathcal{S}_{1}$ $\coloneqq\big\{u_{1}w\colon w\in N_{\mathcal{M}}(e_{\ast})\big\}\cup\big\{u_{2}w\colon w\in N_{\mathcal{M}}(e_{\ast})\big\},$
$\mathcal{S}_{2}$ $\coloneqq\big\{u_{1}w\colon w\in N_{\mathcal{B}}(e_{\ast})\cap V_{1}\big\}\cup\big\{u_{2}w\colon w\in N_{\mathcal{B}}(e_{\ast})\cap V_{1}\big\}$
$\mathcal{S}_{3}$ $\coloneqq\big\{u_{1}w\colon w\in N_{\mathcal{B}}(e_{\ast})\cap V_{3}\big\}\cup\big\{u_{2}w\colon w\in N_{\mathcal{B}}(e_{\ast})\cap V_{3}\big\}.$
Notice that
$|\mathcal{S}_{1}|=2d_{\mathcal{M}}(e_{\ast}),\quad|\mathcal{S}_{2}|+|\mathcal{S}_{3}|=2d_{\mathcal{B}}(e_{\ast}),\quad|\mathcal{S}_{3}|=2d_{\mathcal{B}^{\mathrm{bi}}}(e_{\ast}).$
Let us consider the difference $\lVert\mathcal{G}^{\ast}\rVert_{2}-\lVert\mathcal{G}\rVert_{2}$. Note that for each $e\in\mathcal{S}_{1}$, the codegree of $e$ increased by one, and for each $e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}$, the the codegree of $e$ decreased by one. Therefore,
$\lVert\mathcal{G}^{\ast}\rVert_{2}-\lVert\mathcal{G}\rVert_{2}$ $=\sum_{e\in\binom{V}{2}}\big(d_{\mathcal{G}^{\ast}}^{2}(e)-d_{\mathcal{G}}^{2}(e)\big)$
$=\big(d_{\mathcal{G}^{\ast}}^{2}(e_{\ast})-d_{\mathcal{G}}^{2}(e_{\ast})\big)+\sum_{e\in\mathcal{S}_{1}}\big(d_{\mathcal{G}^{\ast}}^{2}(e)-d_{\mathcal{G}}^{2}(e)\big)+\sum_{e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}}\big(d_{\mathcal{G}^{\ast}}^{2}(e)-d_{\mathcal{G}}^{2}(e)\big)$
$\geq\big(d_{\mathcal{G}^{\ast}}(e_{\ast})-d_{\mathcal{G}}(e_{\ast})\big)\big(d_{\mathcal{G}^{\ast}}(e_{\ast})+d_{\mathcal{G}}(e_{\ast})\big)+\sum_{e\in\mathcal{S}_{1}}2d_{\mathcal{G}}(e)-\sum_{e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}}2d_{\mathcal{G}}(e)$
$=\big(d_{\mathcal{M}}(e_{\ast})-d_{\mathcal{B}}(e_{\ast})\big)\big(d_{\mathcal{G}^{\ast}}(e_{\ast})+d_{\mathcal{G}}(e_{\ast})\big)+\sum_{e\in\mathcal{S}_{1}}2d_{\mathcal{G}}(e)-\sum_{e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}}2d_{\mathcal{G}}(e)$
$\geq-2\xi n^{2}+\sum_{e\in\mathcal{S}_{1}}2d_{\mathcal{G}}(e)-\sum_{e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}}2d_{\mathcal{G}}(e),$ (18)
where the last inequality follows from part (iv) and the trivial upper bound $d_{\mathcal{G}^{\ast}}(e_{\ast})+d_{\mathcal{G}}(e_{\ast})\leq n+n=2n$.
Let $\mathcal{S}_{1}^{h}\coloneqq\left\{e\in\mathcal{S}_{1}\colon d_{\mathcal{M}}(e)\geq\xi^{1/2}n\right\}$. Note that
$d_{\mathcal{M}}(u_{1})+d_{\mathcal{M}}(u_{2})\geq\frac{1}{2}\sum_{e\in\mathcal{S}_{1}^{h}}d_{\mathcal{M}}(e)\geq\frac{1}{2}|\mathcal{S}_{1}^{h}|\cdot\xi^{1/2}n.$
So it follows from $\Delta(\mathcal{M})\leq\xi n^{2}$ that
$|\mathcal{S}_{1}^{h}|\leq\frac{2\big(d_{\mathcal{M}}(u_{1})+d_{\mathcal{M}}(u_{2})\big)}{\xi^{1/2}n}\leq\frac{4\xi n^{2}}{\xi^{1/2}n}=4\xi^{1/2}n.$
Combining this with part (i) and the definition of $\mathcal{S}_{1}^{h}$, we obtain
$\sum_{e\in\mathcal{S}_{1}}d_{\mathcal{G}}(e)\geq\sum_{e\in\mathcal{S}_{1}\setminus\mathcal{S}_{1}^{h}}d_{\mathcal{G}}(e)$ $\geq\sum_{e\in\mathcal{S}_{1}\setminus\mathcal{S}_{1}^{h}}\big(d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(e)-d_{\mathcal{M}}(e)\big)$
$\geq|\mathcal{S}_{1}\setminus\mathcal{S}_{1}^{h}|\left(|V_{1}|+|V_{3}|-1-\xi^{1/2}n\right)$
$\geq\left(|\mathcal{S}_{1}|-4\xi^{1/2}n\right)\left(\frac{2n}{3}-4\xi^{1/2}n\right)$
$\geq\frac{2n|\mathcal{S}_{1}|}{3}-\left(\frac{2n}{3}+|\mathcal{S}_{1}|\right)\cdot 4\xi^{1/2}n>\frac{2n|\mathcal{S}_{1}|}{3}-12\xi^{1/2}n^{2}.$ (19)
Similarly, let $\mathcal{S}_{2}^{h}\coloneqq\left\{e\in\mathcal{S}_{2}\colon d_{\mathcal{B}}(e)\geq\xi^{1/2}n\right\}$. By the same argument as above, the bound $\Delta(\mathcal{B})\leq\xi n^{2}$ implies that $|\mathcal{S}_{2}^{h}|\leq 4\xi^{1/2}n$. It follows from part (v) that $|\mathcal{S}_{3}|\leq 2\xi n$. Therefore,
$\sum_{e\in\mathcal{S}_{2}\cup\mathcal{S}_{3}}d_{\mathcal{G}}(e)$ $=\sum_{e\in\mathcal{S}_{2}\setminus\mathcal{S}_{2}^{h}}d_{\mathcal{G}}(e)+\sum_{e\in\mathcal{S}_{2}^{h}\cup\mathcal{S}_{3}}d_{\mathcal{G}}(e)$
$\leq|\mathcal{S}_{2}|\left(d_{\mathbb{C}[V_{1},V_{2},V_{3}]}+\xi^{1/2}n\right)+\big(|\mathcal{S}_{2}^{h}|+|\mathcal{S}_{3}|\big)n$
$\leq|\mathcal{S}_{2}|\left(\frac{n}{3}+\xi n+\xi^{1/2}n\right)+\big(4\xi^{1/2}n+2\xi n\big)n<\frac{n|\mathcal{S}_{2}|}{3}+10\xi^{1/2}n^{2}.$ (20)
Combining (3.2), (3.2), and (3.2), and applying parts (iii) and (iv), we obtain
$\lVert\mathcal{G}^{\ast}\rVert_{2}-\lVert\mathcal{G}\rVert_{2}$ $\geq-2\xi n^{2}+2\left(\frac{2n|\mathcal{S}_{1}|}{3}-12\xi^{1/2}n^{2}\right)-2\left(\frac{n|\mathcal{S}_{2}|}{3}+10\xi^{1/2}n^{2}\right)$
$\geq\frac{4n}{3}\left(\frac{|\mathcal{S}_{1}|}{2}+\frac{|\mathcal{S}_{1}|}{2}-\frac{|\mathcal{S}_{2}|}{2}\right)-46\xi^{1/2}n^{2}$
$=\frac{4n}{3}\left(d_{\mathcal{M}}(e^{\ast})+d_{\mathcal{M}}(e^{\ast})-d_{\mathcal{B}}(e^{\ast})\right)-46\xi^{1/2}n^{2}$
$\geq\frac{4n}{3}\left(47\xi^{1/2}n-\xi n\right)-46\xi^{1/2}n^{2}>0,$
completing the proof of Lemma 3.6.
Lemma 3.7.
[…]
  1. (i)
    $\max_{i\in[3]}|V_{i}-n/3|\leq\xi n$,
  2. (ii)
    $\max\left\{\Delta(\mathcal{M}),~\Delta(\mathcal{B})\right\}\leq\xi n^{2}$,
  3. (iii)
    $d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})\geq 90\xi^{1/2}n$,
  4. (iv)
    $d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})\geq d_{\mathcal{B}}(e_{\ast})-\xi n$.
Then the new $3$-graph $\mathcal{G}^{\ast}\coloneqq\left(\mathcal{G}\setminus\mathcal{B}(e_{\ast})\right)\cup\mathcal{M}(e_{\ast})$ satisfies
$\lVert\mathcal{G}^{\ast}\rVert_{2}>\lVert\mathcal{G}\rVert_{2}.$
Proof of Lemma 3.7.
Let $\mathcal{G}$, $(V_{1},V_{2},V_{3})$, and $e_{\ast}=\{u_{1},u_{2}\}$ be as in the statement of the lemma. By symmetry, we may assume that $(u_{1},u_{2})\in V_{1}\times V_{2}$. Let
$\mathcal{S}_{1}$ $\coloneqq\big\{u_{1}w\colon w\in N_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})\big\}\cup\big\{u_{2}w\colon w\in N_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})\big\},$
$\mathcal{S}_{2a}$ $\coloneqq\big\{u_{1}w\colon w\in N_{\mathcal{B}}(e_{\ast})\big\}\quad\text{and}\quad\mathcal{S}_{2b}\coloneqq\big\{u_{2}w\colon w\in N_{\mathcal{B}}(e_{\ast})\big\}.$
Let $\mathcal{S}_{2}\coloneqq\mathcal{S}_{2a}\cup\mathcal{S}_{2b}$. Notice that
$|\mathcal{S}_{1}|=2d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast}),\quad|\mathcal{S}_{2a}|=|\mathcal{S}_{2b}|=d_{\mathcal{B}}(e_{\ast}).$
By the same argument as in Lemma 3.6, we have
$\lVert\mathcal{G}^{\ast}\rVert_{2}-\lVert\mathcal{G}\rVert_{2}$ $\geq\big(d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})-d_{\mathcal{B}}(e_{\ast})\big)\big(d_{\mathcal{G}^{\ast}}(e_{\ast})+d_{\mathcal{G}}(e_{\ast})\big)+\sum_{e\in\mathcal{S}_{1}}2d_{\mathcal{G}}(e)-\sum_{e\in\mathcal{S}_{2}}2d_{\mathcal{G}}(e)$
$\geq-2\xi n^{2}+\sum_{e\in\mathcal{S}_{1}}2d_{\mathcal{G}}(e)-\sum_{e\in\mathcal{S}_{2}}2d_{\mathcal{G}}(e).$ (21)
Let $\mathcal{S}_{1}^{h}\coloneqq\left\{e\in\mathcal{S}_{1}\colon d_{\mathcal{M}^{\mathrm{tri}}}(e)\geq\xi^{1/2}n\right\}$ and $\mathcal{S}_{2}^{h}\coloneqq\left\{e\in\mathcal{S}_{2}\colon d_{\mathcal{B}}(e)\geq\xi^{1/2}n\right\}$. Similar to the proof of Lemma 3.6, it follows from part (ii) that $\max\{|\mathcal{S}_{1}^{h}|,~|\mathcal{S}_{2}^{h}|\}\leq 4\xi^{1/2}n$. Similarly, we have
$\sum_{e\in\mathcal{S}_{1}}d_{\mathcal{G}}(e)>\frac{2n|\mathcal{S}_{1}|}{3}-12\xi^{1/2}n^{2},$
and
$\sum_{e\in\mathcal{S}_{2}}d_{\mathcal{G}}(e)$ $=\sum_{e\in\mathcal{S}_{2}\setminus\mathcal{S}_{2}^{h}}d_{\mathcal{G}}(e)+\sum_{e\in\mathcal{S}_{2}^{h}}d_{\mathcal{G}}(e)$
$\leq\sum_{e\in\mathcal{S}_{2}\setminus\mathcal{S}_{2}^{h}}\left(d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(e)+\xi^{1/2}n\right)+|\mathcal{S}_{2}^{h}|\cdot n$
$\leq\sum_{e\in\mathcal{S}_{2a}\setminus\mathcal{S}_{2}^{h}}\left(\frac{2n}{3}+2\xi+\xi^{1/2}n\right)+\sum_{e\in\mathcal{S}_{2b}\setminus\mathcal{S}_{2}^{h}}\left(\frac{n}{3}+\xi+\xi^{1/2}n\right)+4\xi^{1/2}n^{2}$
$\leq\frac{2n|\mathcal{S}_{2a}|}{3}+\frac{n|\mathcal{S}_{2b}|}{3}+9\xi^{1/2}n^{2}.$
Combining these two inequalities with (3.2), and applying parts (iii) and (iv), we obtain
$\lVert\mathcal{G}^{\ast}\rVert_{2}-\lVert\mathcal{G}\rVert_{2}$ $\geq-2\xi n^{2}+2\left(\frac{2n|\mathcal{S}_{1}|}{3}-12\xi^{1/2}n^{2}\right)-2\left(\frac{2n|\mathcal{S}_{2a}|}{3}+\frac{n|\mathcal{S}_{2b}|}{3}+9\xi^{1/2}n^{2}\right)$
$=\frac{4n}{3}\left(\frac{|\mathcal{S}_{1}|}{4}+\frac{|\mathcal{S}_{1}|}{2}-|\mathcal{S}_{2a}|+\frac{|\mathcal{S}_{1}|}{4}-\frac{|\mathcal{S}_{2b}|}{2}\right)-44\xi^{1/2}n^{2}$
$=\frac{4n}{3}\left(\frac{d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})}{2}+\frac{3}{4}\left(d_{\mathcal{M}^{\mathrm{tri}}}(e_{\ast})-d_{\mathcal{B}}(e_{\ast})\right)\right)-44\xi^{1/2}n^{2}$
$\geq\frac{4n}{3}\left(\frac{90\xi^{1/2}n^{2}}{2}-\frac{3\xi n^{2}}{4}\right)-44\xi^{1/2}n^{2}>0.$
This completes the proof of Lemma 3.7.
3.3 Proof of Theorem 1.2
In this subsection, we present the proof of Theorem 1.2.
Proof of Theorem 1.2.
Let $\delta,\delta_{1},\delta_{2},\delta_{3},\delta_{4}>0$ be sufficiently small constants such that $\delta\ll\delta_{1}\ll\delta_{2}\ll\delta_{3}\ll\delta_{4}\ll 1$. Let $n$ be a sufficiently large integer. Let $\mathcal{H}$ be an $n$-vertex $K_{4}^{3}$-free $3$-graph on vertex set $V$ with $\lVert\mathcal{H}\rVert_{2}=\mathrm{ex}_{\ell_{2}}(n,K_{4}^{3})$, that is, with the maximum possible $\ell_{2}$-norm. Let $V_{1}\cup V_{2}\cup V_{3}=V$ be a partition that maximizes the size of the intersection
$\mathcal{H}[V_{1},V_{2},V_{3}]\coloneqq\mathcal{H}\cap\mathbb{C}[V_{1},V_{2},V_{3}].$
Recall that $K[V_{1},V_{2},V_{3}]$ is the complete $3$-partite graph with parts $V_{1}$, $V_{2}$, and $V_{3}$. Also, recall the definitions of $\mathcal{B}$, $\mathcal{M}$, $\mathcal{B}^{\mathrm{int}}$, $\mathcal{B}^{\mathrm{bi}}$, $\mathcal{M}^{\mathrm{tri}}$, and $\mathcal{M}^{\mathrm{bi}}$ from (15), (16), and (17), respectively. We may assume that $\mathcal{B}\neq\emptyset$, since otherwise we are done.
Claim 3.8.
We have $\max\left\{|\mathcal{M}|,~|\mathcal{B}|\right\}\leq 2\delta_{1}n^{3}$.
Proof of Claim 3.8.
Since $\mathbb{C}_{n}$ is $K_{4}^{3}$-free, we have
$\lVert\mathcal{H}\rVert_{2}=\mathrm{ex}_{\ell_{2}}(n,K_{4}^{3})\geq\lVert\mathbb{C}_{n}\rVert_{2}\geq\frac{n^{4}}{6}-\delta n^{4}.$
It follows from Theorem 3.5 that $|\mathcal{H}\triangle\mathbb{C}_{n}|\leq\delta_{1}n^{3}$. Then the choice of the partition $V_{1}\cup V_{2}\cup V_{3}=V$ ensures that $|\mathcal{H}\cap\mathbb{C}[V_{1},V_{2},V_{3}]|\geq|\mathcal{H}\cap\mathbb{C}_{n}|$, which implies
$|\mathcal{H}[V_{1},V_{2},V_{3}]|\geq|\mathbb{C}_{n}|-\delta_{1}n^{3}\geq\frac{5n^{3}}{54}-2\delta_{1}n^{3},$ (22)
It follows that
$|\mathcal{M}|=|\mathbb{C}[V_{1},V_{2},V_{3}]|-|\mathcal{H}[V_{1},V_{2},V_{3}]|\leq|\mathbb{C}_{n}|-|\mathcal{H}[V_{1},V_{2},V_{3}]|\leq 2\delta_{1}n^{3}.$
Note that $|\mathcal{H}\triangle\mathbb{C}_{n}|\leq\delta_{1}n^{3}$ implies $|\mathcal{H}|\leq|\mathbb{C}_{n}|+\delta_{1}n^{3}$. Combining it with (22), we obtain
$|\mathcal{B}|=|\mathcal{H}|-|\mathcal{H}[V_{1},V_{2},V_{3}]|\leq|\mathbb{C}_{n}|+\delta_{1}n^{3}-\left(|\mathbb{C}_{n}|-\delta_{1}n^{3}\right)=2\delta_{1}n^{3}.$
This completes the proof of Claim 3.8.
Let $x_{i}\coloneqq|V_{i}|/n$ for $i\in[3]$.
Claim 3.9.
We have $\max_{i\in[3]}|x_{i}-1/3|\leq 10\delta_{1}^{1/2}$.
Proof of Claim 3.9.
It follows from (22) that
$\frac{5}{54}-2\delta_{1}\leq\frac{|\mathcal{H}[V_{1},V_{2},V_{3}]|}{n^{3}}\leq\frac{|\mathbb{C}[V_{1},V_{2},V_{3}]|}{n^{3}}\leq x_{1}x_{2}x_{3}+\frac{x_{1}^{2}x_{2}}{2}+\frac{x_{2}^{2}x_{3}}{2}+\frac{x_{3}^{2}x_{1}}{2}.$
Combining this with and Lemma 3.3, we obtain
$\max_{i\in[3]}\frac{1}{50}\left(x_{i}-\frac{1}{3}\right)^{2}\leq 2\delta_{1},$
which implies that $\max_{i\in[3]}|x_{i}-1/3|\leq 10\delta_{1}^{1/2}$.
Claim 3.10.
We have $\max\left\{\Delta(\mathcal{B}),~\Delta(\mathcal{M})\right\}\leq 2\delta_{3}n^{2}$.
Proof of Claim 3.10.
Fix an arbitrary vertex $v\in V$. By symmetery, we may assume that $v\in V_{1}$. For convenience, let $L\coloneqq L_{\mathcal{H}}(v)$. Recall from (13) that
$s_{\mathcal{H}}(v)=\lVert L\rVert_{2}+2\sum_{e\in L}d_{\mathcal{H}}(e)-|L|.$ (23)
It follows from Lemma 3.4 that
$s_{\mathcal{H}}(v)\geq s(\mathcal{H})-60n^{2}\geq s(\mathbb{C}_{n})-60n^{2}\geq\frac{2n^{3}}{3}-\delta n^{3}.$ (24)
It follows from the maximality of $\mathcal{H}\cap\mathbb{C}[V_{1},V_{2},V_{3}]$ that moving $v$ from $V_{1}$ to $V_{2}$ and $V_{3}$ will not increase the number of edges in the intersection. Therefore,
$|L[V_{1},V_{2}]|+|L[V_{3}]|$ $\geq|L[V_{1},V_{3}]|+|L[V_{1}]|,$ (25)
$|L[V_{2},V_{3}]|+|L[V_{3}]|$ $\geq|L[V_{1},V_{3}]|+|L[V_{2}]|.$ (26)
Recall that $L[V_{1},V_{2},V_{3}]\coloneqq K[V_{1},V_{2},V_{2}]\cap L$ is the induced $3$-partite subgraph of $L$ with parts $V_{1},V_{2},V_{3}$. For convenience, let $\overline{L}[V_{1},V_{2},V_{3}]\coloneqq L\setminus L[V_{1},V_{2},V_{3}]$. It follows from Claims 3.8 and 3.9 that
$\sum_{e\in L}d_{\mathcal{H}}(e)$ $\leq\sum_{e\in L}\left(d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(e)+d_{\mathcal{B}}(e)\right)$
$=\sum_{e\in\overline{L}[V_{1},V_{2},V_{3}]}d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(e)+\sum_{e\in L[V_{1},V_{2},V_{3}]}d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(e)+\sum_{e\in L}d_{\mathcal{B}}(e)$
$\leq\left(\frac{n}{3}+10\delta_{1}^{1/2}n\right)|L\setminus L[V_{1},V_{2},V_{3}]|+2\left(\frac{n}{3}+10\delta_{1}^{1/2}n\right)|L[V_{1},V_{2},V_{3}]|+3|\mathcal{B}|$
$=\frac{n}{3}\left(|L|+|L[V_{1},V_{2},V_{3}]|\right)+10\delta_{1}^{1/2}n\left(|L|+|L[V_{1},V_{2},V_{3}]|\right)+6\delta_{1}n^{3}$
$\leq\frac{n}{3}\left(|L|+|L[V_{1},V_{2},V_{3}]|\right)+16\delta_{1}^{1/2}n^{3}.$
Combining this with (23) and (24), we obtain
$\lVert L\rVert_{2}+\frac{2n}{3}\left(|L|+|L[V_{1},V_{2},V_{3}]|\right)\geq s_{\mathcal{H}}(v)-16\delta_{1}^{1/2}n^{3}\geq\frac{2}{3}n^{3}-17\delta_{1}^{1/2}n^{3}.$ (27)
Let $\mathcal{T}$ be the set of triangles in $L$ of types $V_{1}V_{2}V_{3}$, $V_{1}V_{1}V_{2}$, $V_{2}V_{2}V_{3}$, and $V_{3}V_{3}V_{1}$, that is,
$\mathcal{T}\coloneqq\big\{S\in\mathbb{C}[V_{1},V_{2},V_{3}]\colon L[S]\cong K_{3}\big\}.$
Suppose, for contradiction, that there exists a triple $S\in\mathcal{T}\setminus\mathcal{M}$. Then the set $S\cup\{v\}$ would induce a copy of $K_{4}^{3}$ in $\mathcal{H}$, contradicting the $K_{4}^{3}$-freeness of $\mathcal{H}$. Thus, $\mathcal{T}\subseteq\mathcal{M}$, and consequently, by Claim 3.8,
$|\mathcal{T}|\leq|\mathcal{M}|\leq 2\delta_{1}n^{3}.$ (28)
Combining this with Claim 3.9, (25), and (26), it follows from Theorem 2.3 that
$\lVert L\rVert_{2}\leq 9\left(\frac{n}{3}\right)^{3}+\delta_{2}n^{3}=\frac{n^{3}}{3}+\delta_{2}n^{3}.$
It then follows from (27) that
$|L[V_{1},V_{2},V_{3}]|+|L|$ $\geq\frac{3}{2n}\left(\frac{2}{3}n^{3}-17\delta_{1}^{1/2}n^{3}-\frac{n^{3}}{3}-\delta_{2}n^{3}\right)\geq\frac{9}{2}\left(\frac{n}{3}\right)^{2}-3\delta_{2}n^{2}.$ (29)
Combining this with (28) and Theorem 2.2, we conclude that for some $i\in[3]$,
$|L[V_{i}]|+|L[V_{i+1}]|+|L[V_{i},V_{i+2}]|\leq\delta_{3}n^{2}.$ (30)
We claim that $i=1$. Suppose to the contrary that $i=2$. Then it follows from (25) that
$|L[V_{1},V_{3}]|+|L[V_{1}]|\leq|L[V_{1},V_{2}]|+|L[V_{3}]|\leq\delta_{3}n^{2}.$
Consequently,
$|L|$ $=\left(|L[V_{2}]|+|L[V_{3}]|+|L[V_{2},V_{1}]|\right)+\left(|L[V_{1},V_{3}]|+|L[V_{1}]|\right)+|L[V_{2},V_{3}]|$
$\leq\delta_{3}n^{2}+\delta_{3}n^{2}+\left(\frac{n}{3}+10\delta_{1}^{1/2}n\right)^{2}<\frac{9}{2}\left(\frac{n}{3}\right)^{2}-3\delta_{2}n^{2},$
a contradiction to (29). Thus, $i\neq 2$. Similarly, using (26), one can show that $i\neq 3$. Therefore, $i=1$, and hence,
$d_{\mathcal{B}}(v)=|L[V_{1}]|+|L[V_{2}]|+|L[V_{1},V_{3}]|\leq\delta_{3}n^{2}.$
This proves that $\Delta(\mathcal{B})\leq\delta_{3}n^{2}$.
It remains to show that $\Delta(\mathcal{M})\leq 2\delta_{3}n^{2}$. It follows from (28) and Theorem 2.11 that
$|L[V_{1},V_{2},V_{3}]|\leq 2\left(\frac{n}{3}\right)^{2}+\delta_{2}n^{2}.$
Combining this with (29), we obtain
$|L|$ $\geq\frac{9}{2}\left(\frac{n}{3}\right)^{2}-3\delta_{2}n^{2}-\left(2\left(\frac{n}{3}\right)^{2}+\delta_{2}n^{2}\right)=\frac{5}{2}\left(\frac{n}{3}\right)^{2}-5\delta_{2}n^{2}.$
It then follows from Claim 3.9 and (30) that
$d_{\mathcal{M}}(v)$ $=d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(v)-\left(|L[V_{1},V_{2}]|+|L[V_{2},V_{3}]|+|L[V_{3}]|\right)$
$\leq d_{\mathbb{C}[V_{1},V_{2},V_{3}]}(v)-\big(|L|-\left(|L[V_{1}]|+|L[V_{2}]|+|L[V_{1},V_{3}]|\right)\big)$
$\leq\frac{5}{2}\left(\frac{n}{3}+10\delta_{1}^{1/2}n\right)^{2}-\left(\frac{5}{2}\left(\frac{n}{3}\right)^{2}-5\delta_{2}n^{2}-\delta_{3}n^{2}\right)\leq 2\delta_{3}n^{2}.$
This proves that $\Delta(\mathcal{M})\leq 2\delta_{3}n^{2}$, and completes the proof of Claim 3.10.
Claim 3.11.
Every bad edge $E\in\mathcal{B}$ contains a pair $e\subseteq E$ such that $d_{\mathcal{M}}(e)\geq n/10$. Moreover, if $E=\{x,y,z\}\in\mathcal{B}^{\mathrm{bi}}$ and $(x,y,z)\in V_{i-1}\times V_{i}\times V_{i}$ for some $i\in[3]$, then
$d_{\mathcal{M}}(yz)\geq\frac{n}{10}\quad\text{or}\quad\max\big\{d_{\mathcal{M}^{\mathrm{tri}}}(xy),~d_{\mathcal{M}^{\mathrm{tri}}}(xz)\big\}\geq\frac{n}{10}.$
Proof of Claim 3.11.
Fix a bad edge $E=\{x,y,z\}\in\mathcal{B}$. First, we consider the case that $E\in\mathcal{B}^{\mathrm{int}}$. By symmetry, we may assume that $E\subseteq V_{1}$. Note that for every $w\in V_{2}$, it follows from the $K_{4}^{3}$-freeness of $\mathcal{H}$ that $\{wxy,wyz,wxz\}\cap\mathcal{M}\neq\emptyset$. Therefore,
$d_{\mathcal{M}}(xy)+d_{\mathcal{M}}(yz)+d_{\mathcal{M}}(xz)\geq|V_{2}|.$
Combining it with Claim 3.10, we obtain
$\max\{d_{\mathcal{M}}(xy),~d_{\mathcal{M}}(yz),~d_{\mathcal{M}}(xz)\}\geq\frac{|V_{2}|}{3}\geq\frac{1}{3}\left(\frac{n}{3}-10\delta_{1}^{1/2}n\right)>\frac{n}{10}.$
Next, we consider the case that $E\in\mathcal{B}^{\mathrm{bi}}$. By symmetry, we may assume that $(x,y,z)\in V_{1}\times V_{2}\times V_{2}$. Note that for every $w\in V_{3}$, it follows from the $K_{4}^{3}$-freeness of $\mathcal{H}$ that $\{wxy,wyz,wxz\}\cap\mathcal{M}\neq\emptyset$. Therefore,
$d_{\mathcal{M}^{\mathrm{tri}}}(xy)+d_{\mathcal{M}^{\mathrm{tri}}}(xz)+d_{\mathcal{M}}(yz)\geq|V_{3}|.$
Combining it with Claim 3.10, we obtain
$\max\{d_{\mathcal{M}^{\mathrm{tri}}}(xy),~d_{\mathcal{M}^{\mathrm{tri}}}(xz),~d_{\mathcal{M}}(yz)\}\geq\frac{|V_{3}|}{3}\geq\frac{1}{3}\left(\frac{n}{3}-10\delta_{1}^{1/2}n\right)>\frac{n}{10}.$
This completes the proof of Claim 3.11.
Recall that $\overline{K}[V_{1},V_{2},V_{3}]=\binom{V_{1}}{2}\cup\binom{V_{2}}{2}\cup\binom{V_{3}}{2}$.
Claim 3.12.
For every $e\in\overline{K}[V_{1},V_{2},V_{3}]$, we have $d_{\mathcal{B}^{\mathrm{bi}}}(e)\leq 3\delta_{3}^{1/2}n$ and, in particular,
$d_{\mathcal{B}}(e)\leq\max_{i\in[3]}\{|V_{i}|\}+3\delta_{3}^{1/2}n\leq\frac{n}{3}+4\delta_{3}^{1/2}n.$
Proof of Claim 3.12.
Suppose to the contrary that there exists $e=\{u_{1},u_{2}\}\in\overline{K}[V_{1},V_{2},V_{3}]$ such that $d_{\mathcal{B}^{\mathrm{bi}}}(e)>3\delta_{3}^{1/2}n$. By symmetry, we may assume that $e\subseteq V_{1}$. Let $N\coloneqq N_{\mathcal{H}}(e)\cap V_{3}$. It follows from the assumption that $|N|=d_{\mathcal{B}^{\mathrm{bi}}}(e)>3\delta_{3}^{1/2}n$. For every pair $\{y,z\}\subseteq N$, it follows from the $K_{4}^{3}$-freeness of $\mathcal{H}$ that $\{u_{1}yz,u_{2}yz\}\cap\mathcal{M}\neq\emptyset$. Therefore,
$\max\{d_{\mathcal{M}}(u_{1}),~d_{\mathcal{M}}(u_{2})\}\geq\frac{1}{2}\left(d_{\mathcal{M}}(u_{1})+d_{\mathcal{M}}(u_{2})\right)\geq\frac{1}{2}\binom{|N|}{2}>2\delta_{3}n^{2},$
a contradiction to Claim 3.10.
Claim 3.13.
For every $e\in\overline{K}[V_{1},V_{2},V_{3}]$, we have $d_{\mathcal{M}}(e)\geq d_{\mathcal{B}}(e)-7\delta_{3}^{1/2}n$ .
Proof of Claim 3.13.
Suppose to the contrary that there exists $e=\{u_{1},u_{2}\}\in\overline{K}[V_{1},V_{2},V_{3}]$ such that this claim fails. By symmetry, we may assume that $e\subseteq V_{1}$.
Let $N_{1}\coloneqq N_{\mathcal{H}}(e)\cap V_{1}$ and $N_{2}\coloneqq N_{\mathcal{H}}(e)\cap V_{2}$. We may assume that $\min\{|N_{1}|,~|N_{2}|\}>2\delta_{3}^{1/2}n$. Indeed, if $|N_{1}|\leq 2\delta_{3}^{1/2}n$, then it follows from Claim 3.12 that
$d_{\mathcal{B}}(e)=d_{\mathcal{B}^{\mathrm{int}}}(e)+d_{\mathcal{B}^{\mathrm{bi}}}(e)=|N_{1}|+d_{\mathcal{B}^{\mathrm{bi}}}(e)\leq 5\delta_{3}^{1/2}n,$
which is clearly at most $d_{\mathcal{M}}(e)+7\delta_{3}^{1/2}n$.
Similarly, if $|N_{2}|\leq 2\delta_{3}^{1/2}n$, then it follows from Claim 3.10 and Claim 3.12 that
$d_{\mathcal{M}}(e)=|V_{2}|-|N_{2}|\geq\frac{n}{3}-10\delta_{1}^{1/2}n-2\delta_{3}^{1/2}n\geq\frac{n}{3}-3\delta_{3}^{1/2}n\geq d_{\mathcal{B}}(e)-7\delta_{3}^{1/2}n.$
Therefore, we may assume that $\min\{|N_{1}|,~|N_{2}|\}>2\delta_{3}^{1/2}n$.
Fix an arbitrary pair $(y,z)\in N_{1}\times N_{2}$. It follows from the $K_{4}^{3}$-freeness of $\mathcal{H}$ that $\{u_{1}yz,u_{2}yz\}\cap\mathcal{M}\neq\emptyset$. Therefore,
$\max\{d_{\mathcal{M}}(u_{1}),~d_{\mathcal{M}}(u_{2})\}$ $\geq\frac{1}{2}\left(d_{\mathcal{M}}(u_{1})+d_{\mathcal{M}}(u_{2})\right)$
$\geq\frac{1}{2}|N_{1}||N_{2}|>\frac{1}{2}\cdot 2\delta_{3}^{1/2}n\cdot 2\delta_{3}^{1/2}n=2\delta_{3}n^{2},$
a contradiction to Claim 3.10.
Define a subfamily of $\mathcal{B}^{\mathrm{bi}}$ as follows:
$\tilde{\mathcal{B}}\coloneqq\big\{\{u,v,w\}\in\mathcal{B}^{\mathrm{bi}}\colon\text{$(u,v,w)\in V_{i}\times V_{i}\times V_{i-1}$ for some $i\in[3]$ and $d_{\mathcal{M}}(uv)\leq\delta_{4}n$}\big\}.$
Claim 3.14.
For every $e\in K[V_{1},V_{2},V_{3}]$, we have $d_{\mathcal{M}^{\mathrm{tri}}}(e)\geq d_{\tilde{\mathcal{B}}}(e)-3\delta_{4}n$.
Proof of Claim 3.14.
Suppose to the contrary that there exists $e=\{u_{1},u_{2}\}\in K[V_{1},V_{2},V_{3}]$ such that this claim fails. By symmetry, we may assume that $(u_{1},u_{2})\in V_{1}\times V_{2}$.
Let $N_{2}\coloneqq N_{\mathcal{H}}(e)\cap V_{2}$ and $N_{3}\coloneqq N_{\mathcal{H}}(e)\cap V_{3}$. We may assume that $\min\{|N_{2}|,~|N_{3}|\}>2\delta_{4}n$. Indeed, if $|N_{2}|\leq 2\delta_{4}n$, then
$d_{\tilde{\mathcal{B}}}(e)\leq d_{\mathcal{B}^{\mathrm{bi}}}(e)=|N_{2}|\leq 2\delta_{4}n\leq d_{\mathcal{M}^{\mathrm{tri}}}(e)+3\delta_{4}n.$
If $|N_{3}|\leq 2\delta_{4}n$, then it follows from Claim 3.10 that
$d_{\mathcal{M}^{\mathrm{tri}}}(e)=|V_{3}|-|N_{3}|$ $\geq\frac{n}{3}-10\delta_{1}^{1/2}n-2\delta_{4}n$
$\geq\frac{n}{3}+10\delta_{1}^{1/2}n-3\delta_{4}n\geq|V_{2}|-3\delta_{4}n\geq d_{\tilde{\mathcal{B}}}(e)-3\delta_{4}n.$
Therefore, we may assume that $\min\{|N_{2}|,~|N_{3}|\}>2\delta_{4}n$.
Fix a vertex $y\in N_{2}$. By the definition of $\tilde{\mathcal{B}}$, we have $d_{\mathcal{M}}(u_{2}y)\leq\delta_{4}<|N_{3}|/2$. Therefore,
$|N_{3}\cap N_{\mathcal{H}}(u_{2}y)|\geq|N_{3}|-d_{\mathcal{M}}(u_{2}y)>\frac{|N_{3}|}{2}.$
For every vertex $z\in N_{3}\cap N_{\mathcal{H}}(u_{2}y)$, it follows from $K_{4}^{3}$-freeness of $\mathcal{H}$ that $u_{1}yz\in\mathcal{M}^{\mathrm{tri}}$. Consequently,
$d_{\mathcal{M}}(u_{1})\geq d_{\mathcal{M}^{\mathrm{tri}}}(u_{1})\geq\frac{1}{2}|N_{2}||N_{3}|>\frac{1}{2}\cdot 2\delta_{4}n\cdot 2\delta_{4}n>2\delta_{3}n,$
a contradiction to Claim 3.10.
We now describe a two-phase procedure that transforms $\mathcal{H}$ into a subgraph of $\mathbb{C}[V_{1},V_{2},V_{3}]$. In Phase One, we remove all bad edges in $\mathcal{B}^{\mathrm{int}}$ and $\mathcal{B}^{\mathrm{bi}}\setminus\tilde{\mathcal{B}}$, while ensuring that the $\ell_{2}$-norm increases (using Lemma 3.6). In Phase Two, we remove all bad edges in $\tilde{\mathcal{B}}\subseteq\mathcal{B}^{\mathrm{bi}}$, again ensuring that the $\ell_{2}$-norm increases (using Lemma 3.7). This yields a contradiction, since it implies that $\lVert\mathcal{H}\rVert_{2}<\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$, contradicting the maximality of $\mathcal{H}$.
Define
$\mathcal{I}$ $\coloneqq\left\{e\in\overline{K}[V_{1},V_{2},V_{3}]\colon d_{\mathcal{M}}(e)\geq\delta_{4}n\right\},\quad\text{and}\quad$
$\mathcal{J}$ $\coloneqq\left\{e\in K[V_{1},V_{2},V_{3}]\colon d_{\mathcal{M}^{\mathrm{tri}}}(e)\geq n/10\right\}.$
We label the elements of $\mathcal{I}$ and $\mathcal{J}$ as
$\mathcal{I}=\{e_{1},\ldots,e_{\ell}\}\quad\text{and}\quad\mathcal{J}=\{\tilde{e}_{1},\ldots,\tilde{e}_{m}\}.$
Phase One: Let $\mathcal{H}_{0}\coloneqq\mathcal{H}$. Suppose that $\mathcal{H}_{i}$ has been defined for some $i\in[0,\ell-1]$. We then let
$\mathcal{H}_{i+1}\coloneqq\big(\mathcal{H}_{i}\setminus\mathcal{B}(e_{i+1})\big)\cup\mathcal{M}(e_{i+1}).$
That is, we remove all bad edges containing $e_{i+1}$ and insert all missing edges containing $e_{i+1}$ (see the figure on the left in Figure 2).
For $i\in[0,\ell]$, let
$\mathcal{M}_{i}\coloneqq\mathbb{C}[V_{1},V_{2},V_{3}]\setminus\mathcal{H}_{i}\quad\text{and}\quad\mathcal{B}_{i}\coloneqq\mathcal{H}_{i}\setminus\mathbb{C}[V_{1},V_{2},V_{3}].$
It follows from the definition of each $\mathcal{H}_{i}$ that
$\mathcal{B}=\mathcal{B}_{0}\supseteq\mathcal{B}_{1}\supseteq\cdots\supseteq\mathcal{B}_{\ell}\quad\text{and}\quad\mathcal{M}=\mathcal{M}_{0}\supseteq\mathcal{M}_{1}\supseteq\cdots\supseteq\mathcal{M}_{\ell}.$
[…]
Claim 3.15.
We have $\lVert\mathcal{H}_{i+1}\rVert_{2}>\lVert\mathcal{H}_{i}\rVert_{2}$ for $i\in[0,\ell-1]$.
Proof of Claim 3.15.
Fix $i\in[0,\ell-1]$. For every $j\in[0,\ell]\setminus\{i\}$, since $\{e_{i},e_{j}\}\subseteq\overline{K}[V_{1},V_{2},V_{3}]$, we have $\mathcal{M}(e_{i})\cap\mathcal{M}(e_{j})=\emptyset$. It follows that
$d_{\mathcal{M}_{i}}(e_{i+1})=d_{\mathcal{M}}(e_{i+1})\geq\delta_{4}n,$ (31)
Moreover, by Claim 3.13,
$d_{\mathcal{M}_{i}}(e_{i+1})=d_{\mathcal{M}}(e_{i+1})\geq d_{\mathcal{B}}(e_{i+1})-7\delta_{3}^{1/2}n\geq d_{\mathcal{B}_{i}}(e_{i+1})-7\delta_{3}^{1/2}n.$ (32)
Applying Lemma 3.6 to $\mathcal{H}_{i}$, with the parameter $\xi$ in that lemma taken to be $7\delta_{3}^{1/2}$, and noting that parts (i), (ii), (iii)(iv), and (iv) are verified by Claim 3.9, Claim 3.10, (31), Claim 3.12, and (32), respectively, we conclude that $\lVert\mathcal{H}_{i+1}\rVert_{2}>\lVert\mathcal{H}_{i}\rVert_{2}$.
Phase Two: Let $\tilde{\mathcal{H}}_{0}\coloneqq\mathcal{H}_{\ell}$, that is, the final $3$-graph obtained at the end of Phase One. Suppose that $\tilde{\mathcal{H}}_{i}$ has been defined for some $i\in[0,m-1]$. We then let
$\tilde{\mathcal{H}}_{i+1}\coloneqq\big(\tilde{\mathcal{H}}_{i}\setminus\mathcal{B}(\tilde{e}_{i})\big)\cup\mathcal{M}^{\mathrm{tri}}(\tilde{e}_{i}).$
For each $i\in[0,m]$, define
$\tilde{\mathcal{B}}_{i}\coloneqq\mathcal{B}_{\tilde{\mathcal{H}}_{i}}[V_{1},V_{2},V_{3}]\quad\text{and}\quad\tilde{\mathcal{M}}_{i}\coloneqq\mathcal{M}_{\tilde{\mathcal{H}}_{i}}^{\mathrm{tri}}[V_{1},V_{2},V_{3}].$
By construction of the sequence $\tilde{\mathcal{H}}_{0},\tilde{\mathcal{H}}_{1},\ldots,\tilde{\mathcal{H}}_{m}$ and the property of $\mathcal{H}_{\ell}$, we have
$\tilde{\mathcal{B}}=\tilde{\mathcal{B}}_{0}\supseteq\tilde{\mathcal{B}}_{1}\supseteq\cdots\supseteq\tilde{\mathcal{B}}_{m}\quad\text{and}\quad\tilde{\mathcal{M}}_{0}\supseteq\tilde{\mathcal{M}}_{1}\supseteq\cdots\supseteq\tilde{\mathcal{M}}_{m}.$
It follows from Claim 3.11 that every bad edge in $\tilde{\mathcal{B}}$ contains an element of $\mathcal{J}$. Thus, the final $3$-graph $\tilde{\mathcal{H}}_{m}$ contains no edge from $\tilde{\mathcal{B}}$, and hence, no bad edge from $\mathcal{B}$. This means that $\tilde{\mathcal{H}}_{m}\subseteq\mathbb{C}[V_{1},V_{2},V_{3}]$.
Claim 3.16.
We have $\Delta(\mathcal{J})\leq 40\delta_{3}n$.
Proof of Claim 3.16.
Suppose to the contrary that this claim fails. Let $v\in V$ be a vertex such that $d_{\mathcal{J}}(v)>40\delta_{3}n$. It follows from the definitions of $\mathcal{J}$ that
$d_{\mathcal{M}}(v)\geq\frac{1}{2}\sum_{u\in N_{\mathcal{J}}(v)}d_{\mathcal{M}^{\mathrm{tri}}}(uv)\geq\frac{d_{\mathcal{J}}(v)}{2}\cdot\frac{n}{10}>2\delta_{3}n^{2},$
a contradiction to Claim 3.10.
Claim 3.17.
For every $i\in[0,m-1]$, the $3$-graph $\tilde{\mathcal{H}}_{i}$ satisfies the following properties:
  1. (i)
    $d_{\tilde{\mathcal{M}}_{i}}(\tilde{e}_{i+1})\geq n/20$,
  2. (ii)
    $d_{\tilde{\mathcal{M}}_{i}}(\tilde{e}_{i+1})\geq d_{\tilde{\mathcal{B}}_{i}}(\tilde{e}_{i+1})-4\delta_{4}n$.
Proof of Claim 3.17.
Fix $i\in[0,m-1]$, and suppose that $\tilde{e}_{i+1}=\{u_{1},u_{2}\}$. By symmetry, we may assume that $(u_{1},u_{2})\in V_{1}\times V_{2}$. Let
$N\coloneqq N_{\tilde{\mathcal{M}}_{i}}(u_{1}u_{2})\subseteq V_{3}\quad\text{and}\quad\hat{N}\coloneqq N_{\mathcal{M}^{\mathrm{tri}}}(u_{1}u_{2})\subseteq V_{3}.$
Since $\tilde{\mathcal{M}}_{i}\subseteq\mathcal{M}^{\mathrm{tri}}$, we have $N\subseteq\hat{N}$. For each $w\in\hat{N}\setminus N$, by the definition of Phase Two, the reason the triple $\{u_{1},u_{2},w\}$ is in $\mathcal{M}^{\mathrm{tri}}$ but not in $\tilde{\mathcal{M}}_{i}$ is that either $\{u_{1},w\}=\tilde{e}_{j}$ or $\{u_{2},w\}=\tilde{e}_{j}$ for some $j\leq i-1$. Hence, by Claim 3.16,
$|\hat{N}\setminus N|\leq d_{\mathcal{J}}(u_{1})+d_{\mathcal{J}}(u_{2})\leq 80\delta_{3}n.$
It follows that
$d_{\tilde{\mathcal{M}}_{i}}(\tilde{e}_{i+1})=|N|\geq|\hat{N}|-80\delta_{3}n=d_{\mathcal{M}^{\mathrm{tri}}}(\tilde{e}_{i+1})-80\delta_{3}n\geq\frac{n}{10}-80\delta_{3}\geq\frac{n}{20},$
which proves part (i).
Combining the inequality above with Claim 3.14, we obtain
$d_{\tilde{\mathcal{M}}_{i}}(\tilde{e}_{i+1})$ $\geq d_{\mathcal{M}^{\mathrm{tri}}}(\tilde{e}_{i+1})-80\delta_{3}n$
$\geq d_{\tilde{\mathcal{B}}}(\tilde{e}_{i+1})-3\delta_{4}n-80\delta_{3}n\geq d_{\tilde{\mathcal{B}}_{i}}(\tilde{e}_{i+1})-4\delta_{4}n,$
which proves (ii).
For $i\in[0,m-1]$, applying Lemma 3.7 to $\tilde{\mathcal{H}}_{i}$, with the parameter $\xi$ in that lemma taken to be $4\delta_{4}$, and noting that parts (i), (ii), (iii), and (iv) are verified by Claim 3.9, Claim 3.10, Claim 3.17 (i), and Claim 3.17 (ii), respectively, we conclude that $\lVert\tilde{\mathcal{H}}_{i+1}\rVert_{2}>\lVert\tilde{\mathcal{H}}_{i}\rVert_{2}$. Consequently, we have the strictly increasing sequence
$\lVert\mathcal{H}\rVert_{2}=\lVert\mathcal{H}_{0}\rVert_{2}<\lVert\mathcal{H}_{1}\rVert_{2}<\cdots<\lVert\mathcal{H}_{\ell}\rVert_{2}=\lVert\tilde{\mathcal{H}}_{0}\rVert_{2}<\cdots<\lVert\tilde{\mathcal{H}}_{m}\rVert_{2}\leq\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert,$
which contradicts the maximality of $\mathcal{H}$. This completes the proof of Theorem 1.2.
4 Concluding remarks
Let $\mathbb{S}_{2}$ denote the unique $3$-graph on $4$ vertices with $2$ edges. Let $\mathrm{N}(\mathbb{S}_{2},\mathcal{H})$ denote the number of copies of $\mathbb{S}_{2}$ in a $3$-graph $\mathcal{H}$. By definition, for every $3$-graph $\mathcal{H}$,
$\lVert\mathcal{H}\rVert_{2}=\sum_{e\in\binom{V(\mathcal{H})}{2}}d_{\mathcal{H}}^{2}(e)=\sum_{e\in\binom{V(\mathcal{H})}{2}}\left(2\tbinom{d_{\mathcal{H}}(e)}{2}+d_{\mathcal{H}}(e)\right)=2\mathrm{N}(\mathbb{S}_{2},\mathcal{H})+3|\mathcal{H}|.$
Thus, Theorem 1.2 implies that, for large $n$, the maximum of $\mathbb{N}(\mathbb{S}_{2},\mathcal{H})+\frac{3}{2}|\mathcal{H}|$ over all $n$-vertex $K_{4}^{3}$-free $3$-graphs $\mathcal{H}$ is uniquely attained by $\mathbb{C}_{n}$. A slight modification of our proof shows that $3/2$ can be replaced by any constant.
Theorem 4.1.
For every real number $\alpha$, there exists $N_{0}=N_{0}(\alpha)$ such that the following holds for every $n\geq N_{0}$. Suppose that $\mathcal{H}$ is an $n$-vertex $K_{4}^{3}$-free $3$-graph. Then
$\mathbb{N}(\mathbb{S}_{2},\mathcal{H})+\alpha|\mathcal{H}|\leq\mathbb{N}(\mathbb{S}_{2},\mathbb{C}_{n})+\alpha|\mathbb{C}_{n}|,$
with equality holds iff $\mathcal{H}\cong\mathbb{C}_{n}$.
In particular, the case $\alpha=0$ establishes exact bound for the generalized Turán number $\mathrm{ex}(n,\mathbb{S}_{2},K_{4}^{3})$ for large $n$.
Given two disjoint sets $V_{1}$ and $V_{2}$, let $\mathbb{B}[V_{1},V_{2}]$ denote the $3$-graph on $V_{1}\cup V_{2}$ consisting of all triples of types $V_{1}V_{1}V_{2}$ and $V_{2}V_{2}V_{1}$ (see Figure 1), that is,
$\mathbb{B}[V_{1},V_{2}]\coloneqq\left\{e\colon\big(|e\cap V_{1}|,|e\cap V_{2}|\big)\in\{(2,1),(1,2)\}\right\}.$
When $V_{1}\cup V_{2}=[n]$ is a balanced partition of $[n]$, we write $\mathbb{B}_{n}$ for $\mathbb{B}[V_{1},V_{2}]$.
The construction $\mathbb{B}_{n}$ does not contain the complete $3$-graph on five vertices $K_{5}^{3}$. It follows that $\pi_{\ell_{2}}(K_{5}^{3})\geq\frac{5}{8}$. Using computer-assisted flag algebra computations, Balogh–Clemen–Lidický [BCL22] proved that $\pi_{\ell_{2}}(K_{5}^{3})\leq\frac{5}{8}$, and thus, $\pi_{\ell_{2}}(K_{5}^{3})=\frac{5}{8}$. They further conjectured (see [BCL22, Conjecture 1.4]) that for large $n$, the construction $\mathbb{B}_{n}$ is the unique extremal construction for the Turán problem of $K_{5}^{3}$ in the $\ell_{2}$-norm. We will confirm this conjecture along with a special case of a related conjecture of Frankl–Gryaznov–Talebanfard [FGT22] in a forthcoming work, after establishing a Turán theorem in vertex-colored graphs.
Acknowledgments
J.H. was supported by the National Key R&D Program of China (No. 2023YFA1010202), the National Natural Science Foundation of China (No. 12071077), and the Central Guidance on Local Science and Technology Development Fund of Fujian Province (No. 2023L3003). X.L. was supported by the Excellent Young Talents Program (Overseas) of the National Natural Science Foundation of China. J.S. was supported by the Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014) and the National Natural Science Foundation of China (No. 12331014).

### v2

@@
 In particular, this implies that the Turán problem for $K_{4}^{3}$ does not exhibit the stability property (see [LM22, Proposition 1.7]), which is likely one reason for the notorious difficulty of the conjecture.
-Successively better upper bounds for $\mathrm{ex}(n,K_{4}^{3})$ have been obtained by many researchers (see e.g., [dC88, CL99, Raz10a, BT11]), with the current record established by the computer-assisted flag algebra framework of Razborov [Raz07].
+Successively better upper bounds for $\mathrm{ex}(n,K_{4}^{3})$ have been obtained by many researchers (see e.g., [dC88, CL99, Raz10, BT11]), with the current record established by the computer-assisted flag algebra framework of Razborov [Raz07].
 $V_{1}$$V_{2}$$V_{3}$$V_{1}$$V_{2}$
@@
 $\pi_{\ell_{2}}(\mathcal{F})\coloneqq\lim_{n\to\infty}\frac{\mathrm{ex}_{\ell_{2}}(n,\mathcal{F})}{\binom{n}{r-1}(n-r+1)^{2}}.$
+Turán problems under the $\ell_{2}$-norm were systematically studied by Balogh–Clemen–Lidický in [BCL22a], and later extended to the more general $(t,p)$-norms as well as to degenerate hypergraphs in [CIL+24, GLMP25].
 The construction $\mathbb{C}_{n}$ shows that $\pi_{\ell_{2}}(K_{4}^{3})\geq\frac{1}{3}$.
-Using Razborov’s flag algebra framework, Balogh–Clemen–Lidický [BCL22] proved, with computer assistance, that $\pi_{\ell_{2}}(K_{4}^{3})\leq\frac{1}{3}$ and thus established the following theorem.
-Theorem 1.1 ([BCL22, Theorem 1.2]).
+Using Razborov’s flag algebra framework, Balogh–Clemen–Lidický [BCL22b] proved, with computer assistance, that $\pi_{\ell_{2}}(K_{4}^{3})\leq\frac{1}{3}$ and thus established the following theorem.
+Theorem 1.1 ([BCL22b, Theorem 1.2]).
 We have $\mathrm{ex}_{\ell_{2}}(n,K_{4}^{3})=\frac{n^{4}}{6}+o(n^{4})$.
-They further conjectured (see [BCL22, Conjecture 1.3]) that for sufficiently large $n$, the construction $\mathbb{C}_{n}$ is the unique extremal configuration for the Turán problem of $K_{4}^{3}$ in the $\ell_{2}$-norm, in contrast to the $\ell_{1}$-norm setting, where exponentially many extremal constructions may exist.
+They further conjectured (see [BCL22b, Conjecture 1.3]) that for sufficiently large $n$, the construction $\mathbb{C}_{n}$ is the unique extremal configuration for the Turán problem of $K_{4}^{3}$ in the $\ell_{2}$-norm, in contrast to the $\ell_{1}$-norm setting, where exponentially many extremal constructions may exist.
 Our main result confirms this conjecture.
@@
 In this section, we present the proof of Theorem 2.3, derived by computer using the flag algebra method of Razborov [Raz07], which is also described in e.g.
-[Raz10b, BaberTalbot11, dCSdOFS16, GGH+22].
+[Raz10, BT11, dCSdOFS16, GGH+22].
 Since this method is well-known by now, we will be very brief.
-In particular, we omit many definitions, referring the reader to [Raz07, Raz10b] for any missing notions.
+In particular, we omit many definitions, referring the reader to [Raz07, Raz10] for any missing notions.
 Roughly speaking, a flag algebra proof using 0-flags on $m$ vertices of an upper bound $u\in\mathbb{R}$ on the given objective function $f$ consists of an identity
@@
 which is asymptotically true for any admissible $\mathcal{H}$ with $|V(\mathcal{H})|\to\infty$, where the $\mathrm{SOS}$-term can be represented as a sum of squares (as described e.g.
-in [Raz10b, Section 3]), each coefficient $c_{F}\in\mathbb{R}$ is non-negative, and $\mathcal{F}_{m}^{0}$ consists of isomorphism types of 0-flags (unlabeled $3$-graphs) with $m$ vertices.
+in [Raz10, Section 3]), each coefficient $c_{F}\in\mathbb{R}$ is non-negative, and $\mathcal{F}_{m}^{0}$ consists of isomorphism types of 0-flags (unlabeled $3$-graphs) with $m$ vertices.
 If $f(\mathcal{H})$ can be represented as a linear combination of the densities of members of $\mathcal{F}_{m}^{0}$ in $\mathcal{H}$ then finding the smallest possible $u$ amounts to solving a semi-definite program (SDP) with $|\mathcal{F}_{m}^{0}|$ linear constraints (so we write the size of $\mathcal{F}_{m}^{0}$ in each case to give the reader some idea of the size of the programs that we had to solve).
@@
 For the remainder of this section, all indices are taken modulo $3$.
-The following statements can be verified by straightforward calculations.
+The following statements can be verified by straightforward calculations (a proof of Fact 3.1 is provided in the appendix of the arXiv version).
 Fact 3.1.
@@
 The following stability theorem of Balogh–Clemen–Lidický will be crucial for our proof.
-Theorem 3.5 ([BCL22, Theorem 1.5]).
+Theorem 3.5 ([BCL22b, Theorem 1.5]).
 […]
@@
 It follows that $\pi_{\ell_{2}}(K_{5}^{3})\geq\frac{5}{8}$.
-Using computer-assisted flag algebra computations, Balogh–Clemen–Lidický [BCL22] proved that $\pi_{\ell_{2}}(K_{5}^{3})\leq\frac{5}{8}$, and thus, $\pi_{\ell_{2}}(K_{5}^{3})=\frac{5}{8}$.
-They further conjectured (see [BCL22, Conjecture 1.4]) that for large $n$, the construction $\mathbb{B}_{n}$ is the unique extremal construction for the Turán problem of $K_{5}^{3}$ in the $\ell_{2}$-norm.
+Using computer-assisted flag algebra computations, Balogh–Clemen–Lidický [BCL22b] proved that $\pi_{\ell_{2}}(K_{5}^{3})\leq\frac{5}{8}$, and thus, $\pi_{\ell_{2}}(K_{5}^{3})=\frac{5}{8}$.
+They further conjectured (see [BCL22b, Conjecture 1.4]) that for large $n$, the construction $\mathbb{B}_{n}$ is the unique extremal construction for the Turán problem of $K_{5}^{3}$ in the $\ell_{2}$-norm.
 We will confirm this conjecture along with a special case of a related conjecture of Frankl–Gryaznov–Talebanfard [FGT22] in a forthcoming work, after establishing a Turán theorem in vertex-colored graphs.
@@
 12331014).
+Appendix
+Proof of Fact 3.1.
+Let $n_{i}\coloneqq|V_{i}|$ for $i\in[3]$.
+It is enough to show that $|n_{i}-n_{j}|\leq 1$ for all $i,j\in[3]$.
+Suppose on the contrary that this is false.
+We will construct a $3$-graph $\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]$ with $V_{1}^{\prime}\cup V_{2}^{\prime}\cup V_{3}^{\prime}=[n]$ such that $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}>\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$, thus proving the lemma.
+Without loss of generality, we may assume that $n_{1}=\max\{n_{1},n_{2},n_{3}\}$.
+Since $n_{1}+n_{2}+n_{3}=n\geq 6$, we have $n_{1}\geq 2$ by average.
+By symmetry, there are two cases.
+Case 1: $n_{1}\geq n_{2}\geq n_{3}$ and $n_{1}\geq n_{3}+2$.
+Let $|V_{1}^{\prime}|=n_{1}-1$, $|V_{2}^{\prime}|=n_{2}$, and $|V_{3}^{\prime}|=n_{3}+1$. Some calculations show that
+$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=$ $n_{1}^{2}n_{3}+2n_{1}(n_{2}^{2}+n_{3}^{2}-n_{2}-n_{3})+n_{2}^{3}-n_{3}^{3}-2n_{2}^{2}n_{3}$
+$-3n_{2}n_{3}^{2}-\frac{7}{2}n_{2}^{2}-\frac{1}{2}n_{3}^{2}+2n_{2}n_{3}+\frac{5}{2}n_{2}+\frac{1}{2}n_{3}.$ (33)
+It suffices to show that this expression is strictly positive.
+When $n_{3}=0$, this is a non-decreasing linear function of $n_{1}$.
+When $n_{3}\geq 1$, this is a quadratic function of $n_{1}$, and it is increasing in $n_{1}\geq 2$.
+So the minimum value of (Appendix) is attained when $n_{1}=\max\{n_{2},n_{3}+2\}$.
+If $n_{1}=n_{2}$, then it follows from (Appendix) that
+$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=3n_{1}^{3}-\frac{11}{2}n_{1}^{2}-n_{3}n_{1}^{2}+\frac{5}{2}n_{1}-n_{3}^{2}n_{1}-n_{3}^{3}-\frac{1}{2}n_{3}^{2}+\frac{1}{2}n_{3},$
+which is increasing in $n_{1}\geq n_{3}+2$.
+So it is enough to assume that $n_{1}=n_{2}=n_{3}+2$.
+Then we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=6n_{3}^{2}+13n_{3}+7>0$.
+If $n_{1}=n_{3}+2$, then $n_{2}$ is one of $n_{3}+2,n_{3}+1$ and $n_{3}$.
+Substituting it into (Appendix), then $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$ is $6n_{3}^{2}+13n_{3}+7$, $6n_{3}^{2}+3n_{3}$ and $6n_{3}^{2}-n_{3}$, respectively.
+The assumptions $n_{1}=n_{3}+2$ and $n\geq 6$ imply that $n_{3}\geq 1$.
+Therefore, we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}>0$, as desired.
+Case 2: $n_{1}\geq n_{3}\geq n_{2}$ and $n_{1}\geq n_{2}+2$.
+Let $|V_{1}^{\prime}|=n_{1}-1$, $|V_{2}^{\prime}|=n_{2}+1$, and $|V_{3}^{\prime}|=n_{3}$. Some calculations show that
+$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=$ $\left(1-n_{1}\right)n_{2}^{2}-2n_{2}\left(n_{1}^{2}+n_{3}^{2}-2n_{1}-n_{3}+1\right)+n_{1}^{3}-\frac{9}{2}n_{1}^{2}$
+$+\frac{13}{2}n_{1}-n_{3}^{3}-\frac{1}{2}n_{3}^{2}+\frac{9}{2}n_{3}+3n_{1}^{2}n_{3}+2n_{1}n_{3}^{2}-8n_{1}n_{3}-3.$ (34)
+This is a quadratic function of $n_{2}$, and it is decreasing in $n_{2}\geq 0$. So it is enough to consider the case $n_{2}=\min\{n_{3},n_{1}-2\}$.
+If $n_{2}=n_{3}$, then it follows from (Appendix) that
+$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\\ =n_{1}^{3}+n_{2}n_{1}^{2}-\frac{9}{2}n_{1}^{2}+n_{2}^{2}n_{1}+\frac{13}{2}n_{1}-4n_{1}n_{2}-3n_{2}^{3}+\frac{5}{2}n_{2}^{2}+\frac{5}{2}n_{2}-3,$ (35)
+which is increasing in $n_{1}\geq n_{2}+2$.
+So it is enough to assume that $n_{1}=n_{2}+2$ and $n_{3}=n_{2}$.
+Then we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=6n_{2}^{2}-n_{2}>0$.
+If $n_{2}=n_{1}-2$, then $n_{3}$ is one of $n_{1}$, $n_{1}-1$ and $n_{1}-2$.
+Substituting it into (Appendix), then $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$ is $6n_{1}^{2}-11n_{1}+5$, $6n_{1}^{2}-15n_{1}+9$ and $6n_{1}^{2}-25n_{1}+26$, respectively.
+The assumptions $n_{1}\geq n_{2}+2$ and $n\geq 6$ imply that $n_{1}\geq 3$.
+Therefore, we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}>0$, as desired.
+Authors
+Levente Bodnár
+Mathematics Institute and DIMAP
+University of Warwick
+Coventry, CV4 7AL, UK
+bodnalev@gmail.com
+Wanfang Chen
+School of Mathematical Sciences
+University of Science and Technology of China
+Hefei, 230026, China
+a372959313@gmail.com
+Jinghua Deng
+Center for Discrete Mathematics
+Fuzhou University
+Fujian, 350003, China
+Jinghua_deng@163.com
+Jianfeng Hou
+Center for Discrete Mathematics
+Fuzhou University
+Fujian, 350003, China
+jfhou@fzu.edu.cn
+Xizhi Liu
+School of Mathematical Sciences
+University of Science and Technology of China
+Hefei, 230026, China
+liuxizhi@ustc.edu.cn
+Jialei Song
+School of Mathematical Sciences
+Key Laboratory of MEA (Ministry of Education) & Shanghai Key Laboratory of PMMP
+East China Normal University
+Shanghai 200241, China
+jialsong@foxmail.com
+Jiabao Yang
+School of Mathematics
+Nanjing University
+Nanjing, 210093, China
+jbyang1215@nju.edu.cn
+Yixiao Zhang
+Center for Discrete Mathematics
+Fuzhou University
+Fujian, 350003, China
+fzuzyx@gmail.com